### v1

# OODTE: A Differential Testing Engine for the ONNX Optimizer

Abstract
With $700$ stars on GitHub and part of the official ONNX repository, the ONNX Optimizer consists of the standard method to apply graph-based optimizations on ONNX models. However, its ability to preserve model accuracy across optimizations, has not been rigorously explored. We propose OODTE, a utility to automatically and thoroughly assess the correctness of the ONNX Optimizer. OODTE follows a simple, yet effective differential testing and evaluation approach that can be easily adopted to other compiler optimizers. In particular, OODTE utilizes a number of ONNX models, then optimizes them and executes both the original and the optimized variants across a user-defined set of inputs, while automatically logging any issues with the optimization process. Finally, for successfully optimized models, OODTE compares the results, and, if any accuracy deviations are observed, it iteratively repeats the process for each pass of the ONNX Optimizer, to localize the root cause of the differences observed. Using OODTE, we sourced well-known $130$ models from the official ONNX Model Hub, used for a wide variety of tasks (classification, object detection, semantic segmentation, text summarization, question and answering, sentiment analysis) from the official ONNX model hub. We observed $9.2$% of all model instances presenting issues leading into the crash of the optimizer, or the generation of an invalid model while using the primary optimizer strategies. In addition, $30$% of the classification models presented accuracy differences across the original and the optimized model variants, while $16.6$% of semantic segmentation and object detection models are also affected, at least to a limited extent. Finally, models related with text analysis and generation were found to be generally robust to the optimization process. Overall, we detected 15 issues, 14 of which were previously unknown, associated with optimizer crashes and accuracy deviations observed in $9$ out of the $47$ available optimizer passes but also the optimizer as a utility in general, which we reported to the developers of the ONNX Optimizer. Our method is simple, yet effective as indicated by the results, and can be easily adapted to other AI model optimization tools and compilers, even beyond the ONNX domain.
I Introduction
Artificial Intelligence has led to a technological revolution, not only within Computer Science but across various aspects of daily life. Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, both researchers and practitioners have developed tools and environments aimed at optimizing DL models [13, 17]. Additionally, widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
Models can also be represented using industry-standard formats, such as ONNX [9]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [5] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question. The ONNX ecosystem also includes additional tools that can further optimize models, potentially at the cost of accuracy—such as the ONNX Quantizer [2]. Developers are typically aware of these trade-offs and apply such tools with consideration of their impact.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the OONX Optimizer Differential Testing Engine (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, and text generation from the official ONNX Model Hub [8]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [6], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. However, our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks. Our methodology builds upon established principles of differential testing [45], which have traditionally been employed for detecting inconsistencies across software variants or transformations.
In total, we observed accuracy deviations in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
Overall, we present the following contributions:
1. OODTE  - a differential testing tool to examine the effects of the ONNX Optimizer to DNNs.
2. A survey of 130 models optimized using the ONNX Optimizer, presenting the effects of the optimizations on model correctness in detail.
3. A set of faults detected using OODTE - reported to the official ONNX Optimizer repository.
II Background
II-A Deep Learning Models
Deep Learning Models (DNNs) are a subset of AI models widely used for prediction tasks across various domains. We focus on the following types of models:
II-A1 Image Classification
One of the earliest and most prevalent applications of DNNs is image classification, typically performed using Convolutional Neural Networks (CNNs). These models employ convolution operations, which function as filters to extract specific features from the input image, using weights and biases, of which values are formed during training. Image classification models are often of small-to-medium size in terms of parameters and layers, taking an image (often preprocessed) as input and outputting a list of top-K prediction pairs (prediction, certainty_score). The sum of the $K$ certainty_score instances equals $1$, with the final operation commonly being a Softmax [47] to amplify the highest prediction probability.
II-A2 Object Detection & Semantic Segmentation
Object detection models extend image classification models by not only classifying objects but also detecting where they are placed within an image, typically using bounding boxes. These models also heavily rely on convolution operations. Object detection is widely used in domains like autonomous driving, where multiple objects need to be detected within a scene. A subset of object detection tasks is semantic segmentation, which classifies each pixel in an image into regions corresponding to different object categories. Each pixel is assigned a specific semantic label. For example, a semantic segmentation model for an autonomous vehicle might label parts of a scene as ”road,” ”cars,” ”pedestrians,” etc., providing a more detailed understanding of the environment than simple object detection.
II-A3 Text Generation
Text generation models are a class of models focused on tasks such as text recognition, summarization, sentiment analysis, and more. These models function similarly to image recognition models, receiving an input and generating an output prediction. In text generation, the model typically receives a sequence of natural language input tokens, generates a prediction for the next token, appends it to the input, and discards the leftmost token of the sequence. This process continues iteratively until a stopping condition is met. This approach is referred to as the sliding window technique.
Text generation models are a subset of Generative AI models, responsible for generating content. A relatively new but rapidly evolving family of models in this category are transformers [54], which use an architecture that processes input data to create internal representations, capturing context. These representations are then used to generate new information, typically in the form of text, by predicting one token at a time sequentially (using the sliding window technique described earlier) until the model output is fully generated.
II-B Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers [12, 32] incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
II-C Comparison Methods
Based on the model type under test, OODTE utilizes a set of well-known and established comparison methods. In particular:
  • For classification models, we use Kendall’s Tau Rank Correlation [28], which measures the correlation between ranked lists. This method is suitable for classification tasks, as it compares the order of ranks, yielding a correlation score in the range of $-1$ (completely negative correlation) to $1$ (completely positive correlation). Kendall’s Tau penalizes order discrepancies between the ranks. In classification tasks, each model typically generates a list of top-K predictions per input. We can then extract these top-K predictions from two model variants and apply Kendall’s Tau to compare their outputs.
  • For Object Detection and Semantic Segmentation models, we utilize the following well-known metrics: (1) Average Intersection-over-Union (IoU) – measures the overlap between predicted and ground truth regions, indicating how well a model recognizes class labels. (2) F1 Score – the harmonic mean of precision and recall, balancing false positives and false negatives. (3) Average Precision (AP) – evaluates how many of the model’s positive predictions were actually correct, focusing on precision across different thresholds. (4) Average Recall (AR) – measures how many actual positives the model correctly identified, focusing on recall across thresholds. (5) Mean Average Precision (mAP) – aggregates the model’s performance across multiple thresholds, evaluating both detection and localization accuracy. To account for varying levels of detection confidence, we performed calculations for metrics (2)-(5) using three different IoU thresholds: $0.5$, $0.75$, and $0.9$. These thresholds are necessary to assess model performance at different levels of localization accuracy, where higher thresholds (e.g., $0.9$) require more precise object detection in the image.
  • For text generation models, we utilize BLEU Score [46], which measures the precision of n-gram matches between the generated text and reference text, evaluating the quality of the generated output. The base refers to the output from the original model, and each optimized model’s output is compared to this base to assess similarity.
  • For binary sentiment analysis, we compare the predictions of each sample between the original and optimized models, calculating the percentage difference in predictions to assess the impact of optimization on sentiment classification.
III System Architecture & Methodology
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator Module, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE  with its different modules is shown in Figure 1. OODTE  has the following intuitive workflow: load models, optimize them, execute both the original and the optimized variants across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
III-A Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models in order to get optimized and run. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [8], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
III-B Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where it only applies a specific set of passes to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
Once a model is optimized with either setting, it then gets propagated to the Runner model, in order to perform inference.
III-C Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model variants. It then performs model inference, using ONNX Runtime [3], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [4] package.
In order to be able to utilize large-size datasets and extract results from heavier models while running on low-end devices, the Runner Module supports execution in-parts: by definition in OODTE configuration, a run can be performed in $C$ chuncks, where each chunk consists of $dataset/C$, with $dataset\%C$ also being the final chunk of the run. This applies both for the original and the optimized model, and the results are extracted and propagated to the Metrics Comparison Module, for evaluation and reporting. Following these approach, we were able to conduct our survey on two devices designed for home use without relying on a cluster for results, described in Section V.
III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [68]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [46] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question and answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. The results can then be visualized, but also they are utilized from OODTE in order to localize faults. As aforementioned, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass and repeats the process, reporting how each model behaves when optimized on each pass in isolation. These results are also stored in separate reports. The whole process is iterative for all chunks of input data, for each optimization (if an optimized model is found to be problematic) and for all models under test.
IV Related Work
Conventional compiler testing is a domain that has been explored [10, 74, 16], involving differential testing [49, 67], as well as including compiler optimizations [11, 27, 19]. However, in terms of testing AI tools and compilers, most of the existing research has focused on (1) the direction of compiler fuzzing and (2) examining context-specific AI tools and APIs, such as DL framework converters.
In terms of compiler fuzzing, HirGen [43] is a utility for Deep Learning compiler fuzzing, focusing on the high-level model intermediate representation (IR) of TVM. TVMFuzz [1] allows generic-purpose TIR fuzzing on Apache TVM, detecting primarily crash and bit manipulation bugs. Tzer [38] is a coverage-guided fuzzer that applies modifications on models compiled in Apache TVM utilizing low-level IR (Tensor IR). GenCoG [70] focuses on generating valid and expressive TVM computation graphs by defining a domain specific language for constraint definition, as well as concolic constraint solving. NNSmith [37] focuses on valid DNN graph generation, usable to examine the compiler for bugs similarly with traditional grammar-based fuzzing techniques (e.g., [74, 16]), and MLIRSmith [69] integrates NNSmith’s fuzzing capabilities in MLIR [32] randomly. Suo et al. [62] also follow an operation dependency approach to guide the fuzzing process in MLIR. DeepDiffer [35] applies low-level IR fuzzing in a coverage-guided manner and comparing results across different DL compilers. SYNTHFUZZ [34] combines grammar-based fuzzing with custom mutation synthesis in order to effectively fuzz MLIR dialects. Li et al. [33] propose a methodology to uncover hardware-related faults by generating fault-sensitive, adversarial tests. Finally, Su et al. [61] fuzz DL compilers in the context of dynamic features, detecting bugs in PyTorch and its underlying tensor compiler.
Regarding testing techniques for context-specific AI tools and APIs, CRADLE [48] performs an execution analysis on the model graph in order to detect faults, while LEMON [71] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. FAME [59] is based on LEMON but enables API mutation and optimizations for layer and weight mutation, guided by the detection of inconcistency faults. Similarly, Audee [20] aims to detect logical, not-a-number bugs, and crashes by applying an exploratory approach in combination with mutation testing. Finally, Louloudakis et al. propose FetaFix [42] - a utility that allows automatic fault detection in model converters, involving ONNX as an intermediate model representation.
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [73] explores the application of differential testing across optimized and unoptimized DNN model variants. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [72] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
To the best of our knowledge, OODTE is the first that thoroughly tests the ONNX Optimizer (and AI optimizers in general), using differential testing in order to detect faults. The closest existing work to our methodology, is DiffChaser [73], DeltaNN [41] and FetaFix [42]. However, these tools considerably differentiate from our contribution. In particular, DiffChaser attempts to detect problems related to AI optimization techniques (such as quantization) related to the models under test, not exploring potential faults introduced from bugs in the optimizer. DeltaNN applies differential testing, but focuses on uncovering behavioral inconsistencies in the models under test when deployed using different configurations, rather than focusing on compiler issues. FetaFix, on the other hand, aims to uncover issues related to converters using differential testing and involving ONNX. Its approach however does not consider model optimizations or AI optimizer tools.
V Experiment Setup
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub. In particular, we selected $93$ models utilized for classification tasks, $30$ models used for object detection and semantic segmentation tasks, and $7$ transformer models [54] utilized for text generation tasks, such as question and answering, (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [51]) and generic-purpose text generation (GPT-2 [50]). We utilize models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - ONNX Runtime v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). However, our approach is generalizable and not restricted to one hub or limited by opset versioning. We selected it because it consists of a standard set of models, refined and prepared for wide use by developers and researchers. The same applies for the models we selected: we focused on models that are mature and strong representatives of their categories (e.g., MobileNet [24], EfficientNet-Lite4 [65] DenseNet121 [25], InceptionV2 [63] RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 56] SSD [39], RoBERTa [40] and GPT-2 [50]), but our methodology can apply to any type of ONNX model. In addition, we utilized text generation models that generated results in a direct, one-off manner (such as RoBERTa), but also in a token retrofitting manner (such as GPT-2). In the latter case, we considered selection of the first, second, and third token prediction, so that our results are as comprehensive as possible.
In terms of computational resources, we ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [3], or an AI compiler such as Apache TVM [12]).
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [57] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text generation, we used SQuAD-v2 [52], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 66], containing $25000$ entries.
To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
  1. 1.
    For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient [28] (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be $<1$ for each input, we counted this input as different.
  2. 2.
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different Intersection-over-Union (IoU) thresholds ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. 3.
    For text generation models performing question answering and summarization, we utilized BLEU score [46] across output texts for each input, monitoring instances where $BLEU\_Score<1$ for each input.
  4. 4.
    For sentiment analysis, we compared each output value consisting of ${0,1}$ across the original and the optimized variants, registering any differences observed.
Finally, we evaluated all existing ONNX Optimizer passes. However, OODTE can also be applied to new or custom passes. Moreover, our methodology can be extended to other optimizers or model optimization APIs (e.g., Apache TVM), provided that the API supports applying passes both individually or as a bundle.
VI Results
Following the execution of the full set of experiments, OODTE successfully identified 15 bugs. Of these, 14 were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
The issues were associated with $9$ of the $47$ passes in the optimizer. However, not all passes are of equal importance; by default, the optimizer utilizes passes related to fuse and eliminate operations. Therefore, $8$ of the discovered issues are deemed more critical, as they are part of the optimizer’s core bundle (default) strategy. Of these, $6$ are directly linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining issues were general and related to the optimizer’s overall usage. We present the impact of these faults both on the optimization process and on the correctness of the optimized models.
TABLE I: Issues detected using OODTE, along with models manifested at. Opset version indicated near each model name.
# Pass Error Type Models
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (11), SSD (12)
2 split_init List index out of range Crash YOLOv3 (10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (7-9), YOLOv2 (10), SSD (12), Inception v1 & v2 (7-9), Tiny-YOLOv3 (11) YOLOv3 (10, 12)
4 split_predict Required inputs are missing. Crash SSD (12), DenseNet121 (7), InceptionV2 (6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (10), DenseNet121 (7) InceptionV2 (6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN R-50-FPN (10)
8 All Passes Models present slightly different results. Diff. See Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (10)
10 eliminate_ consecutive_ idempotent_ops Incompatible Shape Error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (12), R-50-FPN-fp32 (12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN R-50-FPN-int8 (12)
14 All Passes Invalid input scale. Crash VGG-16-BN (7), VGG-19-BN (7), ResNet101_DUC_HDC (7)
15 fuse_bn_into_conv Bad graph structure Diff. See Table II
VI-A Optimizer Crashes
Out of the $15$ issues, $12$ were related to model crashes, $2$ caused model accuracy deviations and $1$ obfuscated the model with redundant initializers, leading to a runtime warning. In particular, we observed that those crashes concerned $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [56], RetinaNet [36], VGG-16, VGG-19 [60], as well as YOLOv3 [55], presented crashes by using the optimizer, in a range of opset instances and in passes such related to the modification of gather nodes, as well as the elimination of idempotent operations. In total, we observed $31$ model instances that either the optimizer crashed upon processing them, or the optimized model was problematic, leading into a runtime crash. However, some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. As a result, from those, $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) ONNX Optimizer passes and are of higer concern.
Overall, classification models were affected towards optimizer crashes to a much lesser extent. However, they were considerably affected in terms of correctness by a subset of faults that did not lead to a crash - which we present below. From text generation models, one issue was related to the crash of the optimizer, and it concerned the rewriting of the input type. However, other text generation models did not present any crash behaviors while optimized using the ONNX optimizer.
VI-B Deviations In Optimized Model Accuracy
Out of the $130$ model variants tested, we observed unexpected accuracy differences between the original and optimized models. This was an unexpected outcome, since ONNX Optimizer transformations are graph-based and do not alter numeric precision, meaning no accuracy deviations were anticipated. Specifically, of the $93$ classification models, $28$ showed accuracy differences, while $5$ of $30$ object detection models had discrepancies. Only one transformer instance (GPT-2) exhibited negligible differences ($<0.001\%$), with all others showing identical results. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the optimizer developers. For instance, ResNet50-caffe2 showed a $12.9\%$ difference in top-10 labels and $2\%$ in top-5 labels, while ShuffleNet presented $0.5\%$ and $1\%$ differences in top-5 and top-10 predictions, respectively.
Figure 2: Illustrative example of problematic label classification and object detection between original (blue) and optimized (orange) model variants, with the bounding box being affected in both cases, with the same (left) and different (right) object label prediction, along with their affected certainty scores.
Figure 3: Model graph malformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
TABLE II: Observed differences across classification predictions for classification (CL) and object detection (OD) models.
Model Type Opset(s) Top 1 Top 5 Top 10 Issue
AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
EfficientNet-Lite4 [65] CL 11 0% 0.04% 1% Graph Malformation
GoogleNet [64] CL 7-9 0% 0% 0.04% Format Version Change
Inception-1 [63] CL 7-9 0% 0% 0.25% Format Version Change
Inception-2 [63] CL 7-9 0% 0.02% 0.05% Format Version Change
MobileNet v2 [24, 58] CL 7-9 0% 0% 0.02% Format Version Change
ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
ShuffleNet-v1 [75] CL 7-9 0% 0.7% 1% Format Version Change
SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
Following runs with OODTE, we concluded that $2$ primary issues were related with accuracy deviations across models:
  • The optimizer automatically upgrades ONNX formats of $v3$ to $v4$, without any warning or indication. This affects models of opsets $7$, $8$, and $9$, which is an outdated opset, but still backwards compatible to the current ONNX Runtime version (v1.21.1), well-represented in the ONNX Model Hub, with $41\%$ of the models in it consisting of any of these opset versions. Of course, not all models within this opset range are affected, yet we found a considerable amount of models presenting differences across their original and optimized variants. We have reported this issue to the optimizer developers and they have acknowledged this behavior for further inspection.
  • The pass fuse_bn_to_conv essentially fuses BatchNormalization and Convolution operators within the model. We observed that this pass was found problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced to the model. In the latter case, the pass attempts to fuse BatchNormalization and Conv nodes, but does so in an incorrect manner, essentially breaking both node properties into chunk nodes. The result can be visually observed in Figure 3. We found that a number of models were producing different accuracies when optimized using this pass. This issue was previously encountered by another user of the optimizer, however no details were given regarding how this issue could be reproduced or what models were affected by it. We raised a comprehensive issue to the ONNX Optimizer developers, providing all the necessary information and details related to its behavior.
Table II showcases the accuracy differences detected in top-K labels for classification and object detection moels, indicating which of the above issues is responsible for this behavior. As an example, Figure 2 illustrates how misclassification and bounding box issues manifest in an input image, with relation to the original (blue) and the optimized (orange) model.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.28% 95.28% 95.28% 86.62% 97.37%
Tiny YOLOv3 [55] 11 99.56% 99.56% 99.56% 90.57% 99.61%
YOLOv2 [53] 9 99.95% 99.95% 99.95% 90.85% 88.50%
YOLOv3 [55] 10 98.83% 98.83% 98.83% 86.66% 98.78%
YOLOv3 [55] 12 98.83% 98.83% 98.83% 86.65% 98.78%
(a) Object Detection Scores with IoU threshold of 50%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [55] 11 99.36% 99.36% 99.36% 89.97% 99.61%
YOLOv2 [53] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [55] 10 97.87% 97.87% 97.87% 85.71% 98.78%
YOLOv3 [55] 12 97.87% 97.87% 97.87% 85.71% 98.78%
(b) Object Detection Scores with IoU threshold of 75%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [55] 11 99.34% 99.34% 99.34% 89.97% 99.61%
YOLOv2 [53] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [55] 10 97.8% 97.8% 97.8% 85.69% 98.78%
YOLOv3 [55] 12 97.8% 97.8% 97.8% 85.69% 98.78%
(c) Object Detection Scores with IoU threshold of 90%.
VI-C Optimizer Warnings
We observed two issues related to the obfuscation of the optimized models with unused initializer entries, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ’INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.
VII Tool Usability
OODTE  can be used directly to test whether the ONNX Optimizer generates optimized models without negatively affecting their accuracy. Specifically, OODTE can verify that the ONNX Optimizer performs correctly on real-scale models of various types and architectures. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
OODTE is an automated utility designed to test the ONNX Optimizer. It can be configured to perform tests both manually and automatically. As a configuration-based tool, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added, whether by the optimizer developers or users implementing custom passes.
In terms of models and datasets, OODTE is adaptable to models provided both locally and fetched from online repositories. By default, OODTE loads models from the official ONNX Model Hub [8], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [7], it can fetch and utilize models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using widely used packages like the datasets [4] package or from local storage. Additionally, OODTE can be easily extended to support any dataset suitable for testing.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and in isolation.
VIII Threats To Validity
In terms of threats to validity, we considered and addressed the following aspects: (1) Comprehension of Experiment Set: We sourced $130$ models from the official ONNX Model Hub, covering classification, object detection, and text generation—categories widely used across various domains. These models are well-established and mature, as the main purpose of the hub is to provide researchers and developers with easily accessible models that require minimal configuration and fine-tuning. The hub offers a broad range of ONNX model versions (opsets), which we considered in our experiments, ranging from version $7$ (the lowest version compatible with the current ONNX Runtime at the time of writing) to version $12$ (the highest available via the ONNX Hub API).
We believe our experiment set is a representative and comprehensive set of models for testing and experimentation. However, OODTE supports models from any hub that complies with ONNX API specification. While our experiment set does not include all models in the hub, we believe it is representative. We excluded models with an opset version lower than $7$ (which are incompatible with the current ONNX and ONNX Runtime version), as well as models for body analysis and super-resolution, since the selected classification and object detection models adequately cover these areas. Additionally, models that only prepare inputs for preprocessing were excluded for being non-generalizable, as were certain transformer models, which were represented by GPT-2.
Overall, we believe our experiment set is comprehensive, representing models commonly used by developers and researchers in the community and industry across a wide range of domains, while adhering to the latest versions supported by ONNX. (2) Optimization Order: We acknowledge that the optimization order can impact model performance [31], a common challenge in compiler research. However, exploring this aspect is beyond the scope of this work. To ensure the validity of our approach, we used two modes in OODTE: (1) the default ONNX Optimizer configuration (which includes the fuse and eliminate passes), and (2) runs where all $47$ optimizer passes were applied atomically. While our approach does not address potential defects arising from the combination of passes, it has been shown to be effective towards detecting correctness issues. (3) Proper Optimization Usage: We followed all guidelines provided by the ONNX Optimizer API to ensure the correct use of each pass. In cases where a pass was expected to produce model graphs that were not strictly compliant with the ONNX format, we took note of this and clearly indicated such instances when reporting the respective issues to the developers. (4) CPU Devices Used: We used both an Intel i5 and an Intel i7 in parallel for our experiments. To ensure consistency across devices, we (a) randomly sampled a subset of models of each type and ran them on both devices, verifying that the results were consistent, and (b) referred to the work of Louloudakis et al. [41], who observed that accuracy remains consistent across different processing devices using optimization. This supported our finding from the sample that results remain consistent across devices.
IX Discussion
We make the following observations from our experiments:
(1) object detection models are more prone to lead into a crash during the optimization process. The vast majority of the crashes we detected were related to object detection and semantic segmentation models. We attribute this to the increased complexity of object detection models in comparison to those utilized only for classification tasks, which are typically simpler in terms of architecture.
(2) The object detection capabilities of problematic models might be considerably affected. In particular:
  • YOLOv3 models show a high average IoU, F1 score, precision, and recall (over 98%). The optimized model misplaces bounding boxes, but not enough to cause misclassifications, as no misclassifications were found in the top-10 label predictions.
  • YOLOv2 showed low IoU values, likely due to its use of multiple bounding boxes per detected object, where small deviations accumulate, negatively affecting the IoU score. However, the high F1 score, average precision, and recall suggest that these deviations do not significantly impact object detection, similarly to how it is illustrated by the left bird instance in Figure 2.
  • In contrast, SSD showed about $4.8$% deviation in F1, precision, and recall between the original and optimized models, leading to $2.9-6$% misclassification rates, despite a high IoU ($>97$%). This is similar to the right bird instance in the illustrative example in Figure 2.
All models presented a considerable mAP score deviation of up to $14.31$%. This indicates that some classes in the model were not properly identified, meaning that the optimizer might affect all object detection and semantic segmentation models towards missing specific objects. This can have a severe effect in models, as despite correctly classifying the objects they detect, they might ”miss” a non-negligible amount objects in the scenes presented to them.
(3) Classification models are more prone to differences in accuracy of the optimized model, compared to the original model. We observed two issues that had an adverse effect towards model classification accuracy. The root causes of such deviations, were attributed to two main issues, (a) one related to implicit, improper change of the ONNX format version for older model opset versions - in which case a large amount of classification models were built using these opsets, and (b) one related to the generation of problematic graph nodes, in relation to convolutional layers (fuse_bn_into_conv pass), a pass that has greater effect in classification models.
(4) Text generation models were affected only to a negligible extent. From the $8$ model instances under test related to text generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model ($4$ when selecting the first token, $2$ when selecting the second and another $2$ when selecting the third token on each prediction). However, we consider this result negligible, as a common practice is to use random token selection across token predictions, and such a difference is so negligible it will essentially not affect model behavior.
Overall, we recognize that the ONNX Optimizer is a tool that performs well for most of the passes and produces good results. We aspire to contribute with this publication, so that the reported issues get fixed by the optimizer developers.
X Future Work
We identify three potential extensions of this work as future research directions:
(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically detecting crashes and accuracy faults. However, given the context of model optimization, it is important to also assess how the optimizer impacts model performance, particularly in terms of execution time (for both CPU and GPU deployments) and other factors such as model size. OODTE can be easily extended to support hardware acceleration model execution (either by using the ONNX Runtime API [3], or by integrating a third-party AI compiler that generates target device code, such as Apache TVM [12]) We leave this as a subject for future work, which could provide valuable insights.
(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models. However, OODTE can be applied to other model types, such as speech recognition, super-resolution, and body recognition. We consider this as a direction for future work.
(3) Examine Optimization Pass Order: OODTE supports optimization with a subset of passes in a defined order. However, our current approach does not account for this, and it can be explored in future contributions.
XI Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected $9.2$% of model instances that crashed when using the primary optimizer strategies, while $30$% of classification models and $16.6$% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
XII Availability
For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.

### v2

@@
 Abstract
-With $700$ stars on GitHub and part of the official ONNX repository, the ONNX Optimizer consists of the standard method to apply graph-based optimizations on ONNX models.
-However, its ability to preserve model accuracy across optimizations, has not been rigorously explored.
-We propose OODTE, a utility to automatically and thoroughly assess the correctness of the ONNX Optimizer.
-OODTE follows a simple, yet effective differential testing and evaluation approach that can be easily adopted to other compiler optimizers.
-In particular, OODTE utilizes a number of ONNX models, then optimizes them and executes both the original and the optimized variants across a user-defined set of inputs, while automatically logging any issues with the optimization process.
-Finally, for successfully optimized models, OODTE compares the results, and, if any accuracy deviations are observed, it iteratively repeats the process for each pass of the ONNX Optimizer, to localize the root cause of the differences observed.
-Using OODTE, we sourced well-known $130$ models from the official ONNX Model Hub, used for a wide variety of tasks (classification, object detection, semantic segmentation, text summarization, question and answering, sentiment analysis) from the official ONNX model hub.
-We observed $9.2$% of all model instances presenting issues leading into the crash of the optimizer, or the generation of an invalid model while using the primary optimizer strategies.
-In addition, $30$% of the classification models presented accuracy differences across the original and the optimized model variants, while $16.6$% of semantic segmentation and object detection models are also affected, at least to a limited extent.
-Finally, models related with text analysis and generation were found to be generally robust to the optimization process.
-Overall, we detected 15 issues, 14 of which were previously unknown, associated with optimizer crashes and accuracy deviations observed in $9$ out of the $47$ available optimizer passes but also the optimizer as a utility in general, which we reported to the developers of the ONNX Optimizer.
-Our method is simple, yet effective as indicated by the results, and can be easily adapted to other AI model optimization tools and compilers, even beyond the ONNX domain.
+With over 700 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models.
+Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated.
+In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer.
+OODTE adopts a straightforward yet powerful differential testing and evaluation methodology, which can be readily adapted for use with other compiler optimizers.
+Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization.
+When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity.
+We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis.
+Our evaluation revealed that $9.2$% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies.
+Additionally, $30$% of classification models and $16.6$% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization.
+In total, OODTE uncovered 15 issues—14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly.
+All findings were reported to the ONNX Optimizer development team.
+OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
 I Introduction
-Artificial Intelligence has led to a technological revolution, not only within Computer Science but across various aspects of daily life.
 Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging.
 As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies.
-To address this, both researchers and practitioners have developed tools and environments aimed at optimizing DL models [13, 17].
-Additionally, widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes.
+To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models [13, 17].
+Widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes for efficient use of hardware.
 These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
@@
 However, the degree to which these optimizations maintain model accuracy remains an open question.
-The ONNX ecosystem also includes additional tools that can further optimize models, potentially at the cost of accuracy—such as the ONNX Quantizer [2].
-Developers are typically aware of these trade-offs and apply such tools with consideration of their impact.
 We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models.
 Ideally, such transformations should not introduce any deviations in model behavior.
-To this end, we propose the OONX Optimizer Differential Testing Engine (OODTE), a framework for performing differential analysis between original and optimized models.
-Using OODTE, we collected 130 models of various types, including classification, object detection, and text generation from the official ONNX Model Hub [8].
+To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a framework for performing differential analysis between original and optimized models.
+Using OODTE, we collected 130 models of various types, including classification, object detection, text comprehension and generation from the official ONNX Model Hub [7].
 Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization.
@@
 We aim to address this gap by introducing a differential testing methodology through OODTE.
-However, our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
-Our methodology builds upon established principles of differential testing [45], which have traditionally been employed for detecting inconsistencies across software variants or transformations.
-In total, we observed accuracy deviations in $34$ out of the $130$ models ($26.1$%).
+Our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
+In total, we observed output differences between the original models and their and optimized counterparts in $34$ out of the $130$ models ($26.1$%).
 Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model.
@@
 II-A Deep Learning Models
-Deep Learning Models (DNNs) are a subset of AI models widely used for prediction tasks across various domains. We focus on the following types of models:
+Deep neural networks (DNNs) are widely used for prediction tasks across domains. We focus on three common categories:
 II-A1 Image Classification
-One of the earliest and most prevalent applications of DNNs is image classification, typically performed using Convolutional Neural Networks (CNNs).
-These models employ convolution operations, which function as filters to extract specific features from the input image, using weights and biases, of which values are formed during training.
-Image classification models are often of small-to-medium size in terms of parameters and layers, taking an image (often preprocessed) as input and outputting a list of top-K prediction pairs (prediction, certainty_score).
-The sum of the $K$ certainty_score instances equals $1$, with the final operation commonly being a Softmax [47] to amplify the highest prediction probability.
-II-A2 Object Detection & Semantic Segmentation
-Object detection models extend image classification models by not only classifying objects but also detecting where they are placed within an image, typically using bounding boxes.
-These models also heavily rely on convolution operations.
-Object detection is widely used in domains like autonomous driving, where multiple objects need to be detected within a scene.
-A subset of object detection tasks is semantic segmentation, which classifies each pixel in an image into regions corresponding to different object categories.
-Each pixel is assigned a specific semantic label.
-For example, a semantic segmentation model for an autonomous vehicle might label parts of a scene as ”road,” ”cars,” ”pedestrians,” etc., providing a more detailed understanding of the environment than simple object detection.
-II-A3 Text Generation
-Text generation models are a class of models focused on tasks such as text recognition, summarization, sentiment analysis, and more.
-These models function similarly to image recognition models, receiving an input and generating an output prediction.
-In text generation, the model typically receives a sequence of natural language input tokens, generates a prediction for the next token, appends it to the input, and discards the leftmost token of the sequence.
-This process continues iteratively until a stopping condition is met.
-This approach is referred to as the sliding window technique.
-Text generation models are a subset of Generative AI models, responsible for generating content.
-A relatively new but rapidly evolving family of models in this category are transformers [54], which use an architecture that processes input data to create internal representations, capturing context.
-These representations are then used to generate new information, typically in the form of text, by predicting one token at a time sequentially (using the sliding window technique described earlier) until the model output is fully generated.
+Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters.
+These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function [46].
+II-A2 Object Detection and Semantic Segmentation
+Object detection extends classification by identifying object locations using bounding boxes.
+Semantic segmentation further refines this by labeling each pixel with a class (e.g., road, pedestrian), offering dense scene understanding.
+Both rely heavily on convolutional operations and are critical in areas like autonomous driving.
+II-A3 Text Comprehension and Generation
+These models handle tasks such as summarization, recognition, and sentiment analysis.
+Text generation models, especially transformers [66], use contextual representations to iteratively predict the next token using a sliding window approach.
+This generative process continues until a stopping condition is met.
 II-B Graph Optimizations
@@
 Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
-II-C Comparison Methods
-Based on the model type under test, OODTE utilizes a set of well-known and established comparison methods. In particular:
-  • For classification models, we use Kendall’s Tau Rank Correlation [28], which measures the correlation between ranked lists.
-This method is suitable for classification tasks, as it compares the order of ranks, yielding a correlation score in the range of $-1$ (completely negative correlation) to $1$ (completely positive correlation).
-Kendall’s Tau penalizes order discrepancies between the ranks.
-In classification tasks, each model typically generates a list of top-K predictions per input.
-We can then extract these top-K predictions from two model variants and apply Kendall’s Tau to compare their outputs.
-  • For Object Detection and Semantic Segmentation models, we utilize the following well-known metrics: (1) Average Intersection-over-Union (IoU) – measures the overlap between predicted and ground truth regions, indicating how well a model recognizes class labels.
-(2) F1 Score – the harmonic mean of precision and recall, balancing false positives and false negatives.
-(3) Average Precision (AP) – evaluates how many of the model’s positive predictions were actually correct, focusing on precision across different thresholds.
-(4) Average Recall (AR) – measures how many actual positives the model correctly identified, focusing on recall across thresholds.
-(5) Mean Average Precision (mAP) – aggregates the model’s performance across multiple thresholds, evaluating both detection and localization accuracy.
-To account for varying levels of detection confidence, we performed calculations for metrics (2)-(5) using three different IoU thresholds: $0.5$, $0.75$, and $0.9$.
-These thresholds are necessary to assess model performance at different levels of localization accuracy, where higher thresholds (e.g., $0.9$) require more precise object detection in the image.
-  • For text generation models, we utilize BLEU Score [46], which measures the precision of n-gram matches between the generated text and reference text, evaluating the quality of the generated output.
-The base refers to the output from the original model, and each optimized model’s output is compared to this base to assess similarity.
-  • For binary sentiment analysis, we compare the predictions of each sample between the original and optimized models, calculating the percentage difference in predictions to assess the impact of optimization on sentiment classification.
+II-C The ONNX Optimizer
+The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes.
+It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency.
+Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead.
+Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing).
+Graph rewrite passes modify graph structure for compatibility, such as rewrite_input_dtype, which adjusts input data types.
+By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
+II-D Model Output Comparison Methods
+Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts.
+In particular:
+Classification models:
+We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation [28], which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
+Object Detection and Semantic Segmentation models:
+Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score.
+For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity.
+For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy.
+These metrics are computed at IoU thresholds of 0.5, 0.75, and 0.9, with the original model as the reference.
+Text generation models:
+We convert token outputs to plain text and use the BLEU Score [45] to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
+Binary sentiment analysis models:
+We measure the percentage of prediction label differences between original and optimized models to evaluate the impact of optimization on classification consistency.
 III System Architecture & Methodology
-Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator Module, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
+Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
 The architecture of OODTE with its different modules is shown in Figure 1.
-OODTE  has the following intuitive workflow: load models, optimize them, execute both the original and the optimized variants across the dataset and compare the results.
+It has the following intuitive workflow: load models, optimize them, execute both the original and the optimized versions across the dataset and compare the results.
 If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model.
@@
 III-A Model Orchestrator Module
-The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models in order to get optimized and run.
+The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution.
 The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository.
@@
 The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart.
-It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where it only applies a specific set of passes to the model.
-It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied, leaving out a number of secondary passes (e.g., ‘rename_input_output‘).
+It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model.
+It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘).
 OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations.
@@
 The Runner model does precisely what its name suggests: executes the models under test.
-In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model variants.
+In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions.
 It then performs model inference, using ONNX Runtime [2], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component.
@@
 It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [3] package.
-In order to be able to utilize large-size datasets and extract results from heavier models while running on low-end devices, the Runner Module supports execution in-parts: by definition in OODTE configuration, a run can be performed in $C$ chuncks, where each chunk consists of $dataset/C$, with $dataset\%C$ also being the final chunk of the run.
-This applies both for the original and the optimized model, and the results are extracted and propagated to the Metrics Comparison Module, for evaluation and reporting.
-Following these approach, we were able to conduct our survey on two devices designed for home use without relying on a cluster for results, described in Section V.
+To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts.
+In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset.
+Any remaining data not evenly divisible is also included in the final chunk of the run.
+This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting.
+By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section V.
 III-D Metrics Comparison Module
@@
 For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [67]), Precision, Recall, and F1 Score [15] given an IoU threshold.
-For text generation models (e.g., GPT-2), it utilizes BLEU [46] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question and answering (Q&A).
+For text generation models (e.g., GPT-2), it utilizes BLEU [45] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A).
 In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
 Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format.
-The results can then be visualized, but also they are utilized from OODTE in order to localize faults.
-As aforementioned, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module.
-Then, the Optimizer Module generates model optimizations per-pass and repeats the process, reporting how each model behaves when optimized on each pass in isolation.
-These results are also stored in separate reports.
-The whole process is iterative for all chunks of input data, for each optimization (if an optimized model is found to be problematic) and for all models under test.
+As mentioned earlier, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module.
+Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.
 IV Related Work
@@
 [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29].
-DiffChaser [73] explores the application of differential testing across optimized and unoptimized DNN model variants.
+DiffChaser [72] explores the application of differential testing across optimized and unoptimized DNN model versions.
 Louloudakis et al.
@@
 V Experiment Setup
-For our experiment set, we sourced $130$ model instances from the ONNX Model Hub.
-In particular, we selected $93$ models utilized for classification tasks, $30$ models used for object detection and semantic segmentation tasks, and $7$ transformer models [54] utilized for text generation tasks, such as question and answering, (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [51]) and generic-purpose text generation (GPT-2 [50]).
-We utilize models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - ONNX Runtime v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API).
-However, our approach is generalizable and not restricted to one hub or limited by opset versioning.
-We selected it because it consists of a standard set of models, refined and prepared for wide use by developers and researchers.
-The same applies for the models we selected: we focused on models that are mature and strong representatives of their categories (e.g., MobileNet [24], EfficientNet-Lite4 [65] DenseNet121 [25], InceptionV2 [63] RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 56] SSD [39], RoBERTa [40] and GPT-2 [50]), but our methodology can apply to any type of ONNX model.
-In addition, we utilized text generation models that generated results in a direct, one-off manner (such as RoBERTa), but also in a token retrofitting manner (such as GPT-2).
-In the latter case, we considered selection of the first, second, and third token prediction, so that our results are as comprehensive as possible.
-In terms of computational resources, we ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning.
+V-1 The Experiment Process
+We evaluated all ONNX Optimizer passes (as detailed in Section II-C).
+For each original model under test, we first applied the default setting of the ONNX Optimizer, which applies all fusion and elimination passes at once, generating the optimized model.
+We then compared the outputs of the original and optimized models according to the model type, as described in Section III-D.
+For optimized models presenting output differences, we performed optimization using only a single optimization pass, resulting in $47$ such single-pass optimized models.
+To identify which pass causes the problematic behavior, we compared the outputs of each single-pass model against those of the original model.
+It is worth noting that OODTE is compatible with new or custom passes as well.
+Furthermore, our methodology generalizes to other optimization frameworks (e.g., Apache TVM), provided they support both bundled and individual pass applications.
+V-2 DNN Models
+For our experiment set, we sourced $130$ model instances from the ONNX Model Hub [7].
+In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet [24], EfficientNet-Lite4 [63] DenseNet121 [25], and InceptionV2 [61]), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 54], and SSD [39]), and $7$ transformer models [66] utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [50]) and generic-purpose text generation (GPT-2 [49]), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks.
+We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API).
+Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
+V-3 Computational Resources
+We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning.
 Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [2], or an AI compiler such as Apache TVM [12]).
+V-4 Datasets
 We selected datasets appropriate to each model task.
 For classification and object detection, we used the ILSVRC2017 [55] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models.
-For text generation, we used SQuAD-v2 [52], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
+For text comprehension and generation, we used SQuAD-v2 [51], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
 Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 64], containing $25000$ entries.
@@
   4. 4.
-    For sentiment analysis, we compared each output value consisting of ${0,1}$ across the original and the optimized variants, registering any differences observed.
-Finally, we evaluated all existing ONNX Optimizer passes.
-However, OODTE can also be applied to new or custom passes.
-Moreover, our methodology can be extended to other optimizers or model optimization APIs (e.g., Apache TVM), provided that the API supports applying passes both individually or as a bundle.
+    For sentiment analysis models, we compared each output value consisting of ${0,1}$ across the original and the optimized versions, registering any differences observed.
 VI Results
-Following the execution of the full set of experiments, OODTE successfully identified 15 bugs.
-Of these, 14 were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps.
+Following the execution of the full set of experiments11footnotemark: 1, OODTE successfully identified $15$ bugs, reported across $11$ separate GitHub issues [9].
+A complete list of these issues, along with the associated models, is shown in Table I.
+Of the 15 bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps.
 We provided comprehensive information and full reproducibility details for this issue.
-The issues were associated with $9$ of the $47$ passes in the optimizer.
-However, not all passes are of equal importance; by default, the optimizer utilizes passes related to fuse and eliminate operations.
-Therefore, $8$ of the discovered issues are deemed more critical, as they are part of the optimizer’s core bundle (default) strategy.
-Of these, $6$ are directly linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining issues were general and related to the optimizer’s overall usage.
-We present the impact of these faults both on the optimization process and on the correctness of the optimized models.
-TABLE I: Issues detected using OODTE, along with models manifested at. Opset version indicated near each model name.
+Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes.
+However, not all passes are equally important.
+By default, the optimizer prioritizes passes related to fuse and eliminate operations.
+As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy.
+Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage.
+We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
+TABLE I: Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (crash or accuracy), and the models in which the error was observed, along with their associated opsets (op.).
 # Pass Error Type Models
-1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (11), SSD (12)
-2 split_init List index out of range Crash YOLOv3 (10, 12)
-3 split_init split_predict Field is required but missing. Crash DenseNet121 (7-9), YOLOv2 (10), SSD (12), Inception v1 & v2 (7-9), Tiny-YOLOv3 (11) YOLOv3 (10, 12)
-4 split_predict Required inputs are missing. Crash SSD (12), DenseNet121 (7), InceptionV2 (6-9)
-5 split_predict Bad spec for node. Crash YOLOv3 (10), DenseNet121 (7) InceptionV2 (6-9)
-6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (12)
-7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN R-50-FPN (10)
-8 All Passes Models present slightly different results. Diff. See Table II
-9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (10)
-10 eliminate_ consecutive_ idempotent_ops Incompatible Shape Error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (12), R-50-FPN-fp32 (12)
-11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN R-50-FPN-qdq (12)
-12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (9)
-13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN R-50-FPN-int8 (12)
-14 All Passes Invalid input scale. Crash VGG-16-BN (7), VGG-19-BN (7), ResNet101_DUC_HDC (7)
-15 fuse_bn_into_conv Bad graph structure Diff. See Table II
+1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
+2 split_init List index out of range. Crash YOLOv3 (op. 10, 12)
+3 split_init split_predict Field is required but missing. Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11) YOLOv3 (op. 10, 12)
+4 split_predict Required inputs are missing. Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
+5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7) InceptionV2 (op. 6-9)
+6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (op. 12)
+7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN: R-50-FPN (op. 10)
+8 All Passes Implicit ONNX format version change. Accuracy Models of cases $1$-$11$ in Table II
+9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (op. 10)
+10 eliminate_ consecutive_ idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
+11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
+12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
+13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN: R-50-FPN-int8 (op. 12)
+14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
+15 fuse_bn_into_conv Bad graph structure. Accuracy Efficient-Lite4 (op. 11), SSD (op. 12)
 VI-A Optimizer Crashes
-Out of the $15$ issues, $12$ were related to model crashes, $2$ caused model accuracy deviations and $1$ obfuscated the model with redundant initializers, leading to a runtime warning.
-In particular, we observed that those crashes concerned $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors.
-In particular, R-CNN models, such as Faster R-CNN R-50-FPN [56], RetinaNet [36], VGG-16, VGG-19 [60], as well as YOLOv3 [55], presented crashes by using the optimizer, in a range of opset instances and in passes such related to the modification of gather nodes, as well as the elimination of idempotent operations.
-In total, we observed $31$ model instances that either the optimizer crashed upon processing them, or the optimized model was problematic, leading into a runtime crash.
-However, some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process.
-As a result, from those, $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) ONNX Optimizer passes and are of higer concern.
-Overall, classification models were affected towards optimizer crashes to a much lesser extent.
-However, they were considerably affected in terms of correctness by a subset of faults that did not lead to a crash - which we present below.
-From text generation models, one issue was related to the crash of the optimizer, and it concerned the rewriting of the input type.
-However, other text generation models did not present any crash behaviors while optimized using the ONNX optimizer.
-VI-B Deviations In Optimized Model Accuracy
-Out of the $130$ model variants tested, we observed unexpected accuracy differences between the original and optimized models.
-This was an unexpected outcome, since ONNX Optimizer transformations are graph-based and do not alter numeric precision, meaning no accuracy deviations were anticipated.
-Specifically, of the $93$ classification models, $28$ showed accuracy differences, while $5$ of $30$ object detection models had discrepancies.
-Only one transformer instance (GPT-2) exhibited negligible differences ($<0.001\%$), with all others showing identical results.
+Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings.
+In particular, we observed crashes were caused by $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes.
+In particular, R-CNN models, such as Faster R-CNN R-50-FPN [54], RetinaNet [36], VGG-16, VGG-19 [58], as well as YOLOv3 [53], presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations.
+In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash.
+Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process.
+Overall, $12$ out of the $31$ model instances ($9.2$%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
+As shown in Table I, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases).
+Overall, object detection models were more prone to optimizer crashes compared to classification models.
+Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults.
+In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2.
+However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
+VI-B Output Label Differences in Optimized Classification & Object Detection Models
+Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models.
+These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks.
+Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected.
+Among the $93$ classification models tested, $28$ ($30.1$%) exhibited discrepancies in output labels.
+Similarly, $5$ out of $30$ ($16.6$%) object detection models presented output differences.
+The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($<0.001\%$) while all others showing identical results across the original and the optimized model versions.
 The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction.
 Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues.
-Our methodology identified these inconsistencies, which have been reported to the optimizer developers.
-For instance, ResNet50-caffe2 showed a $12.9\%$ difference in top-10 labels and $2\%$ in top-5 labels, while ShuffleNet presented $0.5\%$ and $1\%$ differences in top-5 and top-10 predictions, respectively.
-Figure 2: Illustrative example of problematic label classification and object detection between original (blue) and optimized (orange) model variants, with the bounding box being affected in both cases, with the same (left) and different (right) object label prediction, along with their affected certainty scores.
-Figure 3: Model graph malformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
-TABLE II: Observed differences across classification predictions for classification (CL) and object detection (OD) models.
-Model Type Opset(s) Top 1 Top 5 Top 10 Issue
-AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
-DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
-EfficientNet-Lite4 [65] CL 11 0% 0.04% 1% Graph Malformation
-GoogleNet [64] CL 7-9 0% 0% 0.04% Format Version Change
-Inception-1 [63] CL 7-9 0% 0% 0.25% Format Version Change
-Inception-2 [63] CL 7-9 0% 0.02% 0.05% Format Version Change
-MobileNet v2 [24, 58] CL 7-9 0% 0% 0.02% Format Version Change
-ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
-ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
-ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
-ShuffleNet-v1 [75] CL 7-9 0% 0.7% 1% Format Version Change
-SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
-SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
-Following runs with OODTE, we concluded that $2$ primary issues were related with accuracy deviations across models:
-  • The optimizer automatically upgrades ONNX formats of $v3$ to $v4$, without any warning or indication.
-This affects models of opsets $7$, $8$, and $9$, which is an outdated opset, but still backwards compatible to the current ONNX Runtime version (v1.21.1), well-represented in the ONNX Model Hub, with $41\%$ of the models in it consisting of any of these opset versions.
-Of course, not all models within this opset range are affected, yet we found a considerable amount of models presenting differences across their original and optimized variants.
-We have reported this issue to the optimizer developers and they have acknowledged this behavior for further inspection.
-  • The pass fuse_bn_to_conv essentially fuses BatchNormalization and Convolution operators within the model.
-We observed that this pass was found problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced to the model.
-In the latter case, the pass attempts to fuse BatchNormalization and Conv nodes, but does so in an incorrect manner, essentially breaking both node properties into chunk nodes.
-The result can be visually observed in Figure 3.
-We found that a number of models were producing different accuracies when optimized using this pass.
-This issue was previously encountered by another user of the optimizer, however no details were given regarding how this issue could be reproduced or what models were affected by it.
-We raised a comprehensive issue to the ONNX Optimizer developers, providing all the necessary information and details related to its behavior.
-Table II showcases the accuracy differences detected in top-K labels for classification and object detection moels, indicating which of the above issues is responsible for this behavior.
-As an example, Figure 2 illustrates how misclassification and bounding box issues manifest in an input image, with relation to the original (blue) and the optimized (orange) model.
+Our methodology identified these inconsistencies, which have been reported to the ONNX Optimizer developers.
+Figure 2: Illustrative example of problematic label classification and object detection between original unoptimized (blue) and optimized (orange) models.
+The bounding box location is affected for both bird instances, while the object label prediction is affected by model optimization on just the right bird instance, changing the label from bird to jet.
+TABLE II: Output label accuracy differences across classification (CL) and object detection (OD) models, showing the type, opset versions, and accuracy differences for top-1, top-5, and top-10 output predictions, between the original and optimized (using the default ONNX Optimizer setting) versions of each model, along with the issue related to the observed differences.
+# Model Type Opset(s) Top 1 Top 5 Top 10 Issue
+1 AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
+2 DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
+3 GoogleNet [62] CL 7-9 0% 0% 0.04% Format Version Change
+4 Inception-1 [61] CL 7-9 0% 0% 0.25% Format Version Change
+5 Inception-2 [61] CL 7-9 0% 0.02% 0.05% Format Version Change
+6 MobileNet v2 [24, 56] CL 7-9 0% 0% 0.02% Format Version Change
+7 ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
+8 ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
+9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
+10 ShuffleNet-v1 [74] CL 7-9 0% 0.7% 1% Format Version Change
+11 SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
+12 EfficientNet-Lite4 [63] CL 11 0% 0.04% 1% Graph Malformation
+13 SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
+Overall, we detected two primary issues related to output label differences between the original and optimized models:
+Implicit ONNX format version upgrade:
+The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication.
+This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1).
+Despite being outdated, these opset versions are still prevalent, representing approximately 41% of the models in the ONNX Model Hub.
+While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions.
+We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
+Problematic fusion pass:
+The pass fuse_bn_to_conv fuses Batch Normalization and Convolution operators within the model.
+In particular, the pass attempts to fuse BatchNormalization nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node.
+We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model.
+In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula.
+This behavior is illustrated on the right side of Figure 3.
+This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions.
+Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided.
+We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
+Table II shows label differences in top-K outputs for classification and object detection models, highlighting the causes.
+As an example, Figure 2 illustrates how an optimized model (orange) can misclassify and alter bounding boxes compared to the original model (blue).
+Figure 3: Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
+The nodes affected are highlighted.
+The result should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
+TABLE III: Object detection metric differences across the original and optimized versions of object detection models, using the default ONNX Optimizer setting.
+The table presets the name, opset, average F1 score, precision, recall, mean Average Precision (MAP) and average Intersection over Union (IoU scores), for different IoU thresholds ($50$%, $75$%, and $90$%).
 Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
@@
 YOLOv3 [53] 12 98.83% 98.83% 98.83% 86.65% 98.78%
-(a) Object Detection Scores with IoU threshold of 50%.
+(a) IoU threshold of 50%.
 Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
@@
 YOLOv3 [53] 12 97.87% 97.87% 97.87% 85.71% 98.78%
-(b) Object Detection Scores with IoU threshold of 75%.
+(b) IoU threshold of 75%.
 Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
@@
 YOLOv3 [53] 12 97.8% 97.8% 97.8% 85.69% 98.78%
-(c) Object Detection Scores with IoU threshold of 90%.
-VI-C Optimizer Warnings
-We observed two issues related to the obfuscation of the optimized models with unused initializer entries, associated with the eliminate_shape_gather and eliminate_nop_reshape passes.
-When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ’INITIALIZER_ID’.
+(c) IoU threshold of 90%.
+VI-C Bounding Box Differences in Object Detection Models
+We observed notable differences in bounding box outputs between the original and optimized versions of object detection models.
+The results are presented in detail in Table III with the following main findings:
+YOLOv3 Models:
+Both YOLOv3 models (op.set $10$ and $12$) achieve high average IoU, F1 score, precision, and recall—all above $98$%.
+Although the optimized model occasionally misplaces bounding boxes, these errors do not cause misclassifications, as none were found in the top-10 label predictions.
+YOLOv2 Model:
+YOLOv2 shows lower IoU values, likely because it uses multiple bounding boxes per detected object.
+Small differences between the original and optimized models accumulate, lowering IoU.
+However, high F1, precision, and recall indicate that detection performance remains largely unaffected, as demonstrated by the bounding boxes corresponding to the bird on the left in Figure 2.
+SSD Model:
+In contrast, SSD exhibits a $4.8$% drop in F1, precision, and recall between original and optimized models, resulting in 2.92.92.92.96666% misclassification rates despite maintaining a high IoU ($>97$%).
+This behavior is illustrated in Figure 2 by the bounding boxes associated with the bird on the right.
+All models show a substantial mAP difference of up to $14.31$% between original and optimized model versions.
+This suggests the optimizer may cause certain classes to be missed, affecting both object detection and semantic segmentation.
+Consequently, even if detected objects are classified correctly, some objects in the scene might be overlooked, which could significantly impact real-world applications.
+VI-D Optimizer Warnings
+We observed two issues related to the addition of redundant elements, specifically unused initializer entries introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes.
+When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’.
 It is not used by any node and should be removed from the model.
@@
 VII Tool Usability
-OODTE  can be used directly to test whether the ONNX Optimizer generates optimized models without negatively affecting their accuracy.
-Specifically, OODTE can verify that the ONNX Optimizer performs correctly on real-scale models of various types and architectures.
+OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness.
 Currently, OODTE supports differential testing for classification, object detection, and text generation models.
 However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
-OODTE is an automated utility designed to test the ONNX Optimizer.
-It can be configured to perform tests both manually and automatically.
-As a configuration-based tool, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline.
-This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added, whether by the optimizer developers or users implementing custom passes.
-In terms of models and datasets, OODTE is adaptable to models provided both locally and fetched from online repositories.
+In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline.
+This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added.
 By default, OODTE loads models from the official ONNX Model Hub [7], which was used in our experiments.
-However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [7], it can fetch and utilize models without further modification.
+However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [6], it can fetch and use models without further modification.
 The same flexibility applies to datasets.
-OODTE can fetch datasets using widely used packages like the datasets [4] package or from local storage.
-Additionally, OODTE can be easily extended to support any dataset suitable for testing.
-Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and in isolation.
+OODTE can fetch datasets using packages like the datasets [3] package or from local storage.
+Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
 VIII Threats To Validity
-In terms of threats to validity, we considered and addressed the following aspects: (1) Comprehension of Experiment Set: We sourced $130$ models from the official ONNX Model Hub, covering classification, object detection, and text generation—categories widely used across various domains.
-These models are well-established and mature, as the main purpose of the hub is to provide researchers and developers with easily accessible models that require minimal configuration and fine-tuning.
-The hub offers a broad range of ONNX model versions (opsets), which we considered in our experiments, ranging from version $7$ (the lowest version compatible with the current ONNX Runtime at the time of writing) to version $12$ (the highest available via the ONNX Hub API).
-We believe our experiment set is a representative and comprehensive set of models for testing and experimentation.
-However, OODTE supports models from any hub that complies with ONNX API specification.
-While our experiment set does not include all models in the hub, we believe it is representative.
-We excluded models with an opset version lower than $7$ (which are incompatible with the current ONNX and ONNX Runtime version), as well as models for body analysis and super-resolution, since the selected classification and object detection models adequately cover these areas.
-Additionally, models that only prepare inputs for preprocessing were excluded for being non-generalizable, as were certain transformer models, which were represented by GPT-2.
-Overall, we believe our experiment set is comprehensive, representing models commonly used by developers and researchers in the community and industry across a wide range of domains, while adhering to the latest versions supported by ONNX.
-(2) Optimization Order: We acknowledge that the optimization order can impact model performance [31], a common challenge in compiler research.
-However, exploring this aspect is beyond the scope of this work.
-To ensure the validity of our approach, we used two modes in OODTE: (1) the default ONNX Optimizer configuration (which includes the fuse and eliminate passes), and (2) runs where all $47$ optimizer passes were applied atomically.
-While our approach does not address potential defects arising from the combination of passes, it has been shown to be effective towards detecting correctness issues.
-(3) Proper Optimization Usage: We followed all guidelines provided by the ONNX Optimizer API to ensure the correct use of each pass.
-In cases where a pass was expected to produce model graphs that were not strictly compliant with the ONNX format, we took note of this and clearly indicated such instances when reporting the respective issues to the developers.
-(4) CPU Devices Used: We used both an Intel i5 and an Intel i7 in parallel for our experiments.
-To ensure consistency across devices, we (a) randomly sampled a subset of models of each type and ran them on both devices, verifying that the results were consistent, and (b) referred to the work of Louloudakis et al.
-[41], who observed that accuracy remains consistent across different processing devices using optimization.
-This supported our finding from the sample that results remain consistent across devices.
+We list the external and internal threats to validity in our evaluation below.
+VIII-A External Threats
+There are three primary external threats to the validity of our findings.
+(1) Model Selection: Our evaluation included 130 models sourced from the official ONNX Model Hub, covering widely used categories such as classification, object detection, and text comprehension and generation.
+These models were selected for their maturity, accessibility, and broad adoption across academic and industrial domains.
+While we aimed to create a representative and diverse experimental set, the selection is not exhaustive.
+Models with opset versions below 7 were excluded due to incompatibility with the current ONNX and ONNX Runtime versions, which may limit our findings to more recent models.
+Additionally, to keep the evaluation tractable, we did not include certain model types—such as diffusion models and reinforcement learning—which may behave differently under optimization.
+Although our tool supports any ONNX-compliant model hub, the exclusion of these model types presents a potential threat to the generalizability of our results.
+We plan to address this limitation in future work.
+To keep the experiments tractable, we did not evaluate model types such as diffusion models and reinforcement learning in this work; however, we plan to include them in future studies.
+(2) Dataset Selection: Our experimental evaluation relies on widely used, real-world benchmark datasets tailored to the specific tasks of the models under test.
+While this approach ensures relevance and practical applicability—e.g., using ILSVRC [55] for image classification and object detection models like ResNet [22] and GoogleNet [62], and leveraging established HuggingFace datasets [3] for NLP tasks—it may limit generalizability.
+The use of popular datasets could introduce bias toward well-studied scenarios and may not reflect performance on less common or emerging tasks.
+(3) Devices Used: Our experiments were conducted on two hardware platforms: an Intel i5 and an Intel i7 processor.
+While we verified consistency by (a) randomly sampling models of each type and running them on both devices with consistent results, and (b) referencing prior work by Louloudakis et al.
+[41], which found that optimization does not affect accuracy across processing devices, this still presents a potential threat to validity.
+The limited range of hardware may not capture performance or accuracy variations that could arise on other architectures, such as GPUs or edge devices, and may limit the generalizability of our findings.
+VIII-B Internal Threats
+We identify the following internal threat to validity in our findings.
+Optimization Pass Ordering: The order in which optimization passes are applied can significantly affect model performance and behavior, as highlighted in prior work [31].
+This is a well-known challenge in compiler research.
+While this work does not explore the impact of pass ordering in depth, we adopted two evaluation modes in OODTE: (1) the default ONNX Optimizer configuration, which includes commonly used passes such as fuse and eliminate, and (2) a configuration where all 47 available optimization passes were applied atomically.
+Although this strategy does not account for potential issues arising from specific pass combinations or orderings, it has proven effective in uncovering correctness issues in practice.
+Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
 IX Discussion
-We make the following observations from our experiments:
-(1) object detection models are more prone to lead into a crash during the optimization process.
-The vast majority of the crashes we detected were related to object detection and semantic segmentation models.
-We attribute this to the increased complexity of object detection models in comparison to those utilized only for classification tasks, which are typically simpler in terms of architecture.
-(2) The object detection capabilities of problematic models might be considerably affected. In particular:
-  • YOLOv3 models show a high average IoU, F1 score, precision, and recall (over 98%).
-The optimized model misplaces bounding boxes, but not enough to cause misclassifications, as no misclassifications were found in the top-10 label predictions.
-  • YOLOv2 showed low IoU values, likely due to its use of multiple bounding boxes per detected object, where small deviations accumulate, negatively affecting the IoU score.
-However, the high F1 score, average precision, and recall suggest that these deviations do not significantly impact object detection, similarly to how it is illustrated by the left bird instance in Figure 2.
-  • In contrast, SSD showed about $4.8$% deviation in F1, precision, and recall between the original and optimized models, leading to $2.9-6$% misclassification rates, despite a high IoU ($>97$%).
-This is similar to the right bird instance in the illustrative example in Figure 2.
-All models presented a considerable mAP score deviation of up to $14.31$%.
-This indicates that some classes in the model were not properly identified, meaning that the optimizer might affect all object detection and semantic segmentation models towards missing specific objects.
-This can have a severe effect in models, as despite correctly classifying the objects they detect, they might ”miss” a non-negligible amount objects in the scenes presented to them.
-(3) Classification models are more prone to differences in accuracy of the optimized model, compared to the original model.
-We observed two issues that had an adverse effect towards model classification accuracy.
-The root causes of such deviations, were attributed to two main issues, (a) one related to implicit, improper change of the ONNX format version for older model opset versions - in which case a large amount of classification models were built using these opsets, and (b) one related to the generation of problematic graph nodes, in relation to convolutional layers (fuse_bn_into_conv pass), a pass that has greater effect in classification models.
-(4) Text generation models were affected only to a negligible extent.
-From the $8$ model instances under test related to text generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model ($4$ when selecting the first token, $2$ when selecting the second and another $2$ when selecting the third token on each prediction).
-However, we consider this result negligible, as a common practice is to use random token selection across token predictions, and such a difference is so negligible it will essentially not affect model behavior.
-Overall, we recognize that the ONNX Optimizer is a tool that performs well for most of the passes and produces good results.
-We aspire to contribute with this publication, so that the reported issues get fixed by the optimizer developers.
-X Future Work
-We identify three potential extensions of this work as future research directions:
-(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically detecting crashes and accuracy faults.
-However, given the context of model optimization, it is important to also assess how the optimizer impacts model performance, particularly in terms of execution time (for both CPU and GPU deployments) and other factors such as model size.
-OODTE can be easily extended to support hardware acceleration model execution (either by using the ONNX Runtime API [3], or by integrating a third-party AI compiler that generates target device code, such as Apache TVM [12]) We leave this as a subject for future work, which could provide valuable insights.
-(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models.
-However, OODTE can be applied to other model types, such as speech recognition, super-resolution, and body recognition.
-We consider this as a direction for future work.
-(3) Examine Optimization Pass Order: OODTE supports optimization with a subset of passes in a defined order.
-However, our current approach does not account for this, and it can be explored in future contributions.
-XI Conclusion
+We summarize below our findings for the different model types in our experiments.
+(1) Object detection models crash more frequently when optimized.
+Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models.
+This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
+While text generation and comprehension models are also typically large, manual inspection revealed that the ONNX Optimizer applies only a limited set of optimizations to them.
+Most passes are skipped, as many of the layer patterns targeted by the optimizer (e.g., convolutions)—common in classification and object detection models—are not typically present in transformer-based architectures used for language tasks.
+(2) Optimization can reduce the accuracy of object detection models.
+While some object detection models remained relatively stable after optimization, others showed notable performance degradation.
+For example, the YOLOv3 model maintained high detection accuracy with no misclassifications, although it exhibited minor shifts in bounding box positions.
+YOLOv2 also showed stable performance, with slightly lower Intersection over Union (IoU) due to its use of multiple bounding boxes per object, but still achieved strong detection metrics overall.
+In contrast, the SSD model experienced a significant decline in F1 score, precision, and recall, leading to higher misclassification rates despite retaining a high IoU.
+Across all models, we observed substantial drops in mean Average Precision (mAP), with reductions of up to 14.31%, suggesting that optimization may cause certain object classes to be missed.
+This raises concerns about the reliability of optimized models in object detection and segmentation tasks.
+(3) Optimizations can affect label accuracy in classification models.
+We identified two primary causes for differences in label accuracy following optimization.
+(a) The first relates to the implicit and improper upgrade of the ONNX format version in models using older opset versions.
+A significant number of classification models were built on these legacy opsets, making them particularly susceptible to this issue.
+(b) The second involves the generation of problematic graph nodes, particularly in relation to the fuse_bn_into_conv pass.
+This pass, which fuses batch normalization into convolutional layers, has a strong impact on classification models and can lead to unintended changes in model behavior after optimization.
+(4) Outputs from Text comprehension and generation models were not affected by optimization.
+From the $8$ model instances under test related to text comprehension and generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model.
+Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others.
+OODTE  was able to successfully isolate the optimization passes responsible for the output differences for different model types.
+X Conclusion
 We propose OODTE, a tool for the differential testing of the ONNX Optimizer.
@@
 Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
-XII Availability
+XI Availability
 For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.

### v3

@@
 OODTE: A Differential Testing Engine for the ONNX Optimizer
-Abstract
-With over 700 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models.
+Abstract.
+With over 760 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models.
 Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated.
 In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer.
-OODTE adopts a straightforward yet powerful differential testing and evaluation methodology, which can be readily adapted for use with other compiler optimizers.
+OODTE adopts a straightforward yet powerful differential testing, fault localization, and evaluation methodology, which can be readily adapted for use with other compiler optimizers.
 Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization.
@@
 We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis.
-Our evaluation revealed that $9.2$% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies.
-Additionally, $30$% of classification models and $16.6$% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization.
-In total, OODTE uncovered 15 issues—14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly.
+Our evaluation revealed that 9.29.2% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies.
+Additionally, 3030% of classification models and 16.616.6% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization.
+In total, OODTE uncovered 15 issues, 14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly.
 All findings were reported to the ONNX Optimizer development team.
 OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
-I Introduction
+1. Introduction
 Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging.
 As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies.
-To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models [13, 17].
-Widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes for efficient use of hardware.
+To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (speedisallyouneed, ; CompilerGym, ).
+Widely-used AI compilers and frameworks, such as Apache TVM (tvm, ) and MLIR (lattner2020mlircompilerinfrastructureend, ), incorporate their own native sets of optimization passes for efficient use of hardware.
 These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
-Models can also be represented using industry-standard formats, such as ONNX [8].
+Models can also be represented using industry-standard formats, such as ONNX (onnxsite, ).
 By applying transformations to models in these formats, it is possible to reduce model size and improve performance.
-For ONNX in particular, the ONNX Optimizer [4] serves as the default tool for performing graph-level optimizations.
+For ONNX in particular, the ONNX Optimizer (onnxoptimizer, ) serves as the default tool for performing graph-level optimizations.
 As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy.
@@
 Ideally, such transformations should not introduce any deviations in model behavior.
-To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a framework for performing differential analysis between original and optimized models.
-Using OODTE, we collected 130 models of various types, including classification, object detection, text comprehension and generation from the official ONNX Model Hub [7].
-Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization.
-In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
-We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool.
-In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [5], which has received over 4,000 GitHub stars.
+To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a suite that detects and localizes faults in the ONNX Optimizer.
+Problem Being Solved
+OODTE solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations.
+In particular, OODTE employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
+Case Studies
+To evaluate the effectiveness of OODTE, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation.
+All models were obtained from the official ONNX Model Hub (onnxmodelhub, ).
+Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization.
+When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
+We focus on the ONNX Optimizer due to its impact and popularity (over 760 GitHub stars) as a model optimization tool.
+In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (onnxsimplifier, ), which has received over 4,000 GitHub stars.
 However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes.
-We aim to address this gap by introducing a differential testing methodology through OODTE.
-Our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
-In total, we observed output differences between the original models and their and optimized counterparts in $34$ out of the $130$ models ($26.1$%).
-Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model.
-Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation.
-In that case, we were able to provide substantial additional information not previously documented.
-All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
-Overall, we present the following contributions:
-1. OODTE  - a differential testing tool to examine the effects of the ONNX Optimizer to DNNs.
-2. A survey of 130 models optimized using the ONNX Optimizer, presenting the effects of the optimizations on model correctness in detail.
-3. A set of faults detected using OODTE - reported to the official ONNX Optimizer repository.
-II Background
-II-A Deep Learning Models
+With OODTE, we aim to enhance the effectiveness of fault localization and detection in the optimizer.
+However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
+By conducting experiments on our case studies, we observed output discrepancies between the original and optimized models in $34$ of $130$ cases (26.126.1%).
+We also encountered issues with $12$ model instances (9.29.2%) using the primary optimizer passes (fuse and eliminate) where the optimizer crashed or produced invalid models.
+Through analysis, we identified 15 distinct bugs, only one of which had been partially reported before (without a full explanation or reproduction steps).
+For this case, we provided additional insights.
+All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution.
+OODTE effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer.
+These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
+2. Background
+2.1. Deep Learning Models
 Deep Neural Networks (DNNs) are widely used for prediction tasks across domains. We focus on three common categories:
-II-A1 Image Classification
+2.1.1. Image Classification
 Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters.
-These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function [46].
-II-A2 Object Detection and Semantic Segmentation
+These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (softmax, ).
+2.1.2. Object Detection and Semantic Segmentation
 Object detection extends classification by identifying object locations using bounding boxes.
@@
 Both rely heavily on convolutional operations and are critical in areas like autonomous driving.
-II-A3 Text Comprehension and Generation
+2.1.3. Text Comprehension and Generation
 These models handle tasks such as summarization, recognition, and sentiment analysis.
-Text generation models, especially transformers [66], use contextual representations to iteratively predict the next token using a sliding window approach.
+Text generation models, especially transformers (attention, ), use contextual representations to iteratively predict the next token using a sliding window approach.
 This generative process continues until a stopping condition is met.
-II-B Graph Optimizations
+2.2. Graph Optimizations
 For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance.
@@
 This may involve techniques such as node pruning, node fusion, and constant folding, among others.
-Similar to conventional software, these optimizations are applied to DNNs, with AI compilers [12, 32] incorporating such techniques.
+Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (tvm, ; lattner2020mlircompilerinfrastructureend, ) incorporating such techniques.
 Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
-II-C The ONNX Optimizer
+2.3. The ONNX Optimizer
 The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes.
@@
 Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing).
-Graph rewrite passes modify graph structure for compatibility, such as rewrite_input_dtype, which adjusts input data types.
+Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types.
+In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_pre-
+dict
+, which isolate the initializer and inference subgraphs, respectively.
 By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
-II-D Model Output Comparison Methods
+2.4. Model Output Comparison Methods
 Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts.
 In particular:
-Classification models:
-We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation [28], which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
-Object Detection and Semantic Segmentation models:
-Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score.
+Classification models::
+We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (kendall, ), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
+Object Detection & Semantic Segmentation models::
+Ea-
+ch prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score.
 For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity.
@@
 These metrics are computed at IoU thresholds of $0.5$, $0.75$, and $0.9$, with the original model as the reference.
-Text generation models:
-We convert token outputs to plain text and use the BLEU Score [45] to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
-Binary sentiment analysis models:
+Text generation models::
+We convert token outputs to plain text and use the BLEU Score (BLEU, ) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
+Binary sentiment analysis models::
 We measure the percentage of prediction label differences between original and optimized models to evaluate the impact of optimization on classification consistency.
-III System Architecture & Methodology
-Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
-The architecture of OODTE with its different modules is shown in Figure 1.
-It has the following intuitive workflow: load models, optimize them, execute both the original and the optimized versions across the dataset and compare the results.
-If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model.
-This is done to localize the optimization pass responsible for causing an output difference.
-All results are stored in report files, both for reporting and localization purposes.
-In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
-OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from.
-This is managed by the Model Orchestrator Module.
-The modules within OODTE  in Figure 1 is explained in the following Sections.
-III-A Model Orchestrator Module
+3. System Architecture & Methodology
+Figure 1. Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
+The architecture of OODTE and its modules is shown in Figure 1.
+Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results.
+If output discrepancies are detected, OODTE iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference.
+All results are stored in report files for analysis and fault localization.
+OODTE consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
+It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module.
+The modules in Figure 1 are detailed in the following sections.
+3.1. Model Orchestrator Module
 The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution.
-The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository.
-At the moment, OODTE supports the official ONNX Model Hub [7], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories.
+The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository.
+At the moment, OODTE supports the official ONNX Model Hub (onnxmodelhub, ), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories.
 In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification.
 Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
-III-B Optimizer Module
+3.2. Optimizer Module
 The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart.
-It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model.
+It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model, applied in a user-defined order specified in the system configuration.
 It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘).
 OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations.
-This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth.
-OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
-Once a model is optimized with either setting, it then gets propagated to the Runner model, in order to perform inference.
-III-C Runner Module
+This feature is essential throughout the OODTE pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth.
+To identify the pass responsible for problematic behavior, OODTE first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified.
+Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
+3.3. Runner Module
 The Runner model does precisely what its name suggests: executes the models under test.
 In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions.
-It then performs model inference, using ONNX Runtime [2], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component.
+It then performs model inference, using ONNX Runtime (onnxruntime, ), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component.
 OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews.
-It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [3] package.
+It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (datasets, ) package.
 To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts.
@@
 This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting.
-By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section V.
-III-D Metrics Comparison Module
+By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section 5.
+3.4. Metrics Comparison Module
 Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module.
 This module performs comparison against the results, by selecting the appropriate comparator based on the model type.
-For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28].
-For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [67]), Precision, Recall, and F1 Score [15] given an IoU threshold.
-For text generation models (e.g., GPT-2), it utilizes BLEU [45] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A).
+For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (kendall, ).
+For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (IoUBoundingBox, )), Precision, Recall, and F1 Score (F1Score, ) given an IoU threshold.
+For text generation models (e.g., GPT-2), it utilizes BLEU (BLEU, ) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A).
 In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
+3.5. Fault Localization
 Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format.
-As mentioned earlier, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module.
+If a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module.
 Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.
-IV Related Work
-Conventional compiler testing is a domain that has been explored [10, 73, 16], involving differential testing [48, 65], as well as including compiler optimizations [11, 27, 19].
-However, in terms of testing AI tools and compilers, most of the existing research has focused on (1) the direction of compiler fuzzing and (2) examining context-specific AI tools and APIs, such as DL framework converters.
-In terms of compiler fuzzing, HirGen [43] is a utility for Deep Learning compiler fuzzing, focusing on the high-level model intermediate representation (IR) of TVM.
-TVMFuzz [1] allows generic-purpose TIR fuzzing on Apache TVM, detecting primarily crash and bit manipulation bugs.
-Tzer [38] is a coverage-guided fuzzer that applies modifications on models compiled in Apache TVM utilizing low-level IR (Tensor IR).
-GenCoG [69] focuses on generating valid and expressive TVM computation graphs by defining a domain specific language for constraint definition, as well as concolic constraint solving.
-NNSmith [37] focuses on valid DNN graph generation, usable to examine the compiler for bugs similarly with traditional grammar-based fuzzing techniques (e.g., [73, 16]), and MLIRSmith [68] integrates NNSmith’s fuzzing capabilities in MLIR [32] randomly.
-Suo et al.
-[60] also follow an operation dependency approach to guide the fuzzing process in MLIR.
-DeepDiffer [35] applies low-level IR fuzzing in a coverage-guided manner and comparing results across different DL compilers.
-SYNTHFUZZ [34] combines grammar-based fuzzing with custom mutation synthesis in order to effectively fuzz MLIR dialects.
-Li et al.
-[33] propose a methodology to uncover hardware-related faults by generating fault-sensitive, adversarial tests.
-Finally, Su et al.
-[59] fuzz DL compilers in the context of dynamic features, detecting bugs in PyTorch and its underlying tensor compiler.
-Regarding testing techniques for context-specific AI tools and APIs, CRADLE [47] performs an execution analysis on the model graph in order to detect faults, while LEMON [70] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model.
-FAME [57] is based on LEMON but enables API mutation and optimizations for layer and weight mutation, guided by the detection of inconcistency faults.
-Similarly, Audee [20] aims to detect logical, not-a-number bugs, and crashes by applying an exploratory approach in combination with mutation testing.
-Finally, Louloudakis et al.
-propose FetaFix [42] - a utility that allows automatic fault detection in model converters, involving ONNX as an intermediate model representation.
-Overall, limited work has been done in the direction of testing AI compiler optimizations.
-Choi et al.
-[14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29].
-DiffChaser [72] explores the application of differential testing across optimized and unoptimized DNN model versions.
-Louloudakis et al.
-propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations.
-Xiao et al.
-[71] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
-To the best of our knowledge, OODTE is the first that thoroughly tests the ONNX Optimizer (and AI optimizers in general), using differential testing in order to detect faults.
-The closest existing work to our methodology, is DiffChaser [72], DeltaNN [41] and FetaFix [42].
-However, these tools considerably differentiate from our contribution.
-In particular, DiffChaser attempts to detect problems related to AI optimization techniques (such as quantization) related to the models under test, not exploring potential faults introduced from bugs in the optimizer.
-DeltaNN applies differential testing, but focuses on uncovering behavioral inconsistencies in the models under test when deployed using different configurations, rather than focusing on compiler issues.
-FetaFix, on the other hand, aims to uncover issues related to converters using differential testing and involving ONNX.
-Its approach however does not consider model optimizations or AI optimizer tools.
-V Experiment Setup
-V-1 The Experiment Process
-We evaluated all ONNX Optimizer passes (as detailed in Section II-C).
-For each original model under test, we first applied the default setting of the ONNX Optimizer, which applies all fusion and elimination passes at once, generating the optimized model.
-We then compared the outputs of the original and optimized models according to the model type, as described in Section III-D.
-For optimized models presenting output differences, we performed optimization using only a single optimization pass, resulting in $47$ such single-pass optimized models.
-To identify which pass causes the problematic behavior, we compared the outputs of each single-pass model against those of the original model.
-It is worth noting that OODTE is compatible with new or custom passes as well.
-Furthermore, our methodology generalizes to other optimization frameworks (e.g., Apache TVM), provided they support both bundled and individual pass applications.
-V-2 DNN Models
-For our experiment set, we sourced $130$ model instances from the ONNX Model Hub [7].
-In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet [24], EfficientNet-Lite4 [63] DenseNet121 [25], and InceptionV2 [61]), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 54], and SSD [39]), and $7$ transformer models [66] utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [50]) and generic-purpose text generation (GPT-2 [49]), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks.
-We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API).
+OODTE applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model.
+In addition, OODTE registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model).
+Upon detecting an issue or a problematic behavior, OODTE attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model.
+This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
+4. Related Work
+In this work, we compare our work with contributions related to AI compilers, focusing on the following key aspects: (1) the ability to localize faults associated with graph optimization passes, (2) the detection of compiler bugs affecting The types of bugs considered are the ones resulting in (a) model accuracy (inference results), (b) compiler crashes, (c) successful compilations, which, however, produce malformed or incorrect models that fail to run (e.g., models with incorrect input nodes or parameters), and (d) issues impacting execution times, (3) the capability to localize faults in the ONNX domain, particularly within ONNX Optimizer, and (4) the detection of faults in graph optimization passes.
+Table 1 summarizes the features supported by each contribution.
+For each tool, Table 1 shows: (1) the underlying technique, (2) support for bug detection in ONNX, (3) the ability to localize faults (Flt Loc.), (4) the types of bugs detected and the corresponding graph optimization passes involved (Gr.
+Opt.
+Passes), (5) detection of label accuracy faults (Acc.), (6) detection of crash faults (Cr.), (7) detection of malformed models produced despite successful compilation (Malf.), (8) detection of bugs affecting execution times (Ex.
+T.), and (9) the compilers supported/evaluated.
+Table 1.
+Comparison of features between OODTE and related work.
+Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt.
+Loc.), bug types, graph optimization passes (Gr.
+Opt.
+Passes), detection of accuracy changes (Acc.), crashes (Cr.), generation of malformed models due to miscompilation (Malf.), and issues affecting model execution times(T.) bugs, as well as the compilers explored on each contribution.
+:*: Limited support of feature, $\dagger:$ Does not consider ONNX Optimizer or compile-time optimizations associated with ONNX, $\ddagger$}: Uses ONNX as an intermediate representation, but does not perform fault detection associated to it.
+Tool Technique ON NX? Flt. Loc. Bug Types Gr. Opt. Passes Acc. Cr. Malf. T. Compilers
+OODTE
+Per-pass fault
+localization,
+differential
+testing
+Logic, API versioning,
+graph input/parameters/
+transformations
+Fuse/
+eliminate/
+rewrite
+ONNX
+Optimizer (onnxoptimizer, )
+NNSmith (NNSmith, )
+Grammar-
+based fuzzing
+†,‡ $\times$
+Conversions, semantics,
+gen.transformations,
+graph rewrites
+Conv2D
+inputs
+rewrite
+$\times$ $\times$
+TVM (tvm, ),
+PyTorch JIT (pytorch, ),
+TensorRT (tensorrt, ),
+PT2 (pytorch2, ), ONNX
+Runtime (onnxruntime, ),
+TF XLA (tfxla, ),
+TFLite (tensorflow, )
+MLIRSmith (MLIRSmith, )
+Grammar-
+based fuzzing
+$\times$ $\times$
+Conversions (dialects),
+generic transformations,
+logic (dialects)
+$\times$ $\times$ $\times$ $\times$
+MLIR (MLIR, ),
+LLVM (LLVM, )
+MLIRod (SuoMLIR, )
+Operator
+dependency
+analysis/
+fuzzing
+$\times$ $\times$
+Conversions (dialects),
+gen. transformations
+$\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
+SYNTH-
+FUZZ (synthfuzz, )
+Grammar-
+based fuzzing
+$\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
+HirGen (HirGen, )
+Graph IR
+fuzzing
+$\times$ $\times$
+Numeric computations,
+exceptions, types, memory
+$\times$ $\times$ $\times$ $\times$ TVM (tvm, )
+TVMFuzz (TVMFuzz, )
+Low-level
+IR fuzzing
+$\times$ $\times$
+Numeric computations,
+graph rewrites
+Constant
+folding
+$\times$ $\times$ TVM (tvm, )
+Tzer (TZER, )
+Coverage-
+guided fuzzing
+(Low-level IR)
+$\times$ $\times$
+API inconsistency/misuse,
+memory, types
+$\times$ $\times$ TVM (tvm, )
+GenCoG (GenCoG, )
+Constraint
+solving,
+High-level
+IR Fuzzing
+$\times$ $\times$
+Memory, tensor shape
+/data, operator attributes,
+logic
+Operator
+fusion
+$\times$ $\times$ TVM (tvm, )
+DeepDiffer (DeepDiffer, )
+Low-level IR/
+coverage-
+guided fuzzing,
+differential
+testing
+$\times$ $\times$
+Type, memory, API misuse,
+numeric computations,
+exception handling
+$\times$ $\times$ $\times$ TVM (tvm, )
+WhiteFox (WhiteFox, )
+LM analysis,
+fuzzing
+$\times$ $\times$
+Memory, API misuse,
+concurrency, logic,
+security vulnerabilities
+$\times$ $\times$
+TF XLA (tfxla, ),
+PyTorch
+Inductor (pytorch, ),
+LLVM (LLVM, )
+TorchProbe (Su_2023, )
+Feature-
+based fuzzing
+$\times$ $\times$
+Memory, API misuse,
+logic
+$\times$ $\times$ $\times$
+PyTorch (pytorch, ),
+Triton (TilletTriton, )
+MT-
+DLComp (Xiao_2022, )
+Metamorphic
+testing
+$\times$ $\times$
+Memory, dead code,
+layout optimization,
+numeric accuracy
+Conv2D
+layout
+change
+$\times$ $\times$
+TVM (tvm, ),
+Glow (Glow, ),
+NNFusion (NNFusion, ),
+TF XLA (tfxla, )
+FetaFix (FetaFix, )
+Fault localiza-
+tion,
+differential
+testing
+$\times$
+Preprocessing, input
+dimensions, parameters,
+hyperparameters, tensor
+shape, graph structure
+$\times$ $\times$ $\times$ $\times$
+Converters across
+PyTorch (pytorch, ),
+Keras (chollet2015keras, ),
+TF (tensorflow, ), TFLite (tensorflow, )
+In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (DeepDiffer, )) or LLM analysis (e.g., WhiteFox (WhiteFox, )).
+Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization.
+Among these works, only FetaFix (FetaFix, ) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process.
+In contrast, OODTE systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
+In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs.
+For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (NNSmith, )) to expose inconsistencies in model inference.
+In contrast, for label inconsistency bugs that OODTE detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (ILSVRC17, ), IMDB (imdbsentiment, )) and model architectures.
+Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (TVMFuzz, ; WhiteFox, ).
+The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (TZER, ; DeepDiffer, ) using empirical measurements to evaluate performance-related faults.
+OODTE currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries), while the integration of precise execution time measurements is left as subject for future work.
+Regarding support for ONNX in Table 1, OODTE is the only tool capable of localizing faults in the ONNX Optimizer.
+Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (WhiteFox, ), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
+Existing work on bug detection within graph optimization passes (shown as Gr.
+Opt.
+Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations.
+In contrast, OODTE addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
+5. Experiment Setup
+5.1. The Experiment Process
+We evaluated all ONNX Optimizer passes (Section 2.3).
+For each model, we first applied the default setting, which runs all fusion and elimination passes, and compared the outputs of the original and optimized models as described in Section 3.4.
+When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue.
+OODTE also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
+5.2. DNN Models
+For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (onnxmodelhub, ).
+In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet (mobilenet, ), EfficientNet-Lite4 (efficientnet, ) DenseNet121 (densenet, ), and InceptionV2 (inception, )), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (retinanet, ), Mask/Faster R-CNN R-50-FPN (mask-rcnn, ; faster-rcnn, ), and SSD (ssd, )), and $7$ transformer models (attention, ) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (BERT, )), sentiment analysis (RoBERTa (liu2020roberta, )), text summarization (T5 (T5Paper, )) and generic-purpose text generation (GPT-2 (gpt2, )), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks.
+We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API).
 Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
-V-3 Computational Resources
+5.3. Computational Resources
 We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning.
-Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [2], or an AI compiler such as Apache TVM [12]).
-V-4 Datasets
+Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (onnxruntime, ), or an AI compiler such as Apache TVM (tvm, )).
+5.4. Datasets
 We selected datasets appropriate to each model task.
-For classification and object detection, we used the ILSVRC2017 [55] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models.
-For text comprehension and generation, we used SQuAD-v2 [51], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
-Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 64], containing $25000$ entries.
+For classification and object detection, we used the ILSVRC2017 (ILSVRC17, ) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (resnet, ) and BERT (BERT, )).
+For text comprehension and generation, we used SQuAD-v2 (rajpurkar2018knowdontknowunanswerable, ), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
+Additionally, for summarization and sentiment analysis, we used the IMDB dataset (moviedataset, ; imdbsentiment, ), containing $25000$ entries.
 To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
-  1. 1.
-    For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient [28] (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$.
-If the KT value was found to be $<1$ for each input, we counted this input as different.
-  2. 2.
-    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different Intersection-over-Union (IoU) thresholds ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall.
+  1. (1)
+    For image recognition models performing classification tasks, we utilized the Kendall Rank Correlation Coefficient (kendall, ) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$.
+If the KT value was found to be <1<1 for each input, we counted this input as different.
+  2. (2)
+    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall.
 We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
-  3. 3.
-    For text generation models performing question answering and summarization, we utilized BLEU score [45] across output texts for each input, monitoring instances where $BLEU\_Score<1$ for each input.
-  4. 4.
-    For sentiment analysis models, we compared each output value consisting of ${0,1}$ across the original and the optimized versions, registering any differences observed.
-VI Results
-Following the execution of the full set of experiments11footnotemark: 1, OODTE successfully identified $15$ bugs, reported across $11$ separate GitHub issues [9].
-A complete list of these issues, along with the associated models, is shown in Table I.
+  3. (3)
+    For text generation models performing question answering and summarization, we utilized BLEU score (BLEU, ) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
+  4. (4)
+    For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
+6. Results
+Following the execution of the full set of experiments.
+A complete list of these issues, along with the associated models, is shown in Table 2.
+The developers have acknowledged the issues we reported, and we have provided them with full reproducibility details for each reported bug.
 Of the $15$ bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps.
@@
 As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy.
-Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage.
+Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage.
 We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
-TABLE I: Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (crash or accuracy), and the models in which the error was observed, along with their associated opsets (op.).
-# Pass Error Type Models
-1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
-2 split_init List index out of range. Crash YOLOv3 (op. 10, 12)
-3 split_init split_predict Field is required but missing. Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11) YOLOv3 (op. 10, 12)
-4 split_predict Required inputs are missing. Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
-5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7) InceptionV2 (op. 6-9)
-6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (op. 12)
-7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN: R-50-FPN (op. 10)
-8 All Passes Implicit ONNX format version change. Accuracy Models of cases $1$-$11$ in Table II
-9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (op. 10)
-10 eliminate_ consecutive_ idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
-11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
-12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
-13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN: R-50-FPN-int8 (op. 12)
-14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
-15 fuse_bn_into_conv Bad graph structure. Accuracy Efficient-Lite4 (op. 11), SSD (op. 12)
-VI-A Optimizer Crashes
+6.1. Optimizer Crashes
 Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings.
-In particular, we observed crashes were caused by $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes.
-In particular, R-CNN models, such as Faster R-CNN R-50-FPN [54], RetinaNet [36], VGG-16, VGG-19 [58], as well as YOLOv3 [53], presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations.
+In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes.
+In particular, R-CNN models, such as Faster R-CNN R-50-FPN (faster-rcnn, ), RetinaNet (retinanet, ), VGG-16, VGG-19 (vgg, ), as well as YOLOv3 (Redmon2018YOLOv3AI, ), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations.
 In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash.
 Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process.
-Overall, $12$ out of the $31$ model instances ($9.2$%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
-As shown in Table I, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases).
+Overall, $12$ out of the $31$ model instances (9.29.2%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
+As shown in Table 2, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases).
 Overall, object detection models were more prone to optimizer crashes compared to classification models.
@@
 However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
-VI-B Output Label Differences in Optimized Classification & Object Detection Models
+Table 2.
+Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
+# Pass Error Type Models
+1 lift_lexical_references
+Removes used
+value references.
+Crash
+EfficientNet-Lite4 (op. 11), SSD (op. 12)
+2 split_init
+List index
+out of range.
+Crash YOLOv3 (op. 10, 12)
+3
+split_init
+split_predict
+Required Field
+is missing.
+Crash
+DenseNet121 (op. 7-9), YOLOv2 (op. 10),
+SSD (op. 12), Tiny-YOLOv3 (op. 11),
+YOLOv3 (op. 10, 12)
+4 split_predict
+Required inputs
+are missing.
+Crash
+SSD (op. 12), DenseNet121 (op. 7),
+InceptionV2 (op. 6-9)
+5 split_predict Bad spec for node. Crash
+YOLOv3 (op. 10), DenseNet121 (op. 7),
+InceptionV2 (op. 6-9)
+6 split_predict
+Mismatch in number
+of input elements
+Crash
+FCN ResNet-50-int8 (op. 12)
+7 eliminate_shape_gather
+Indices value
+not equal to
+node dims size
+Crash
+Faster R-CNN: R-50-FPN (op. 10)
+8 All Passes
+Implicit ONNX
+format version
+update.
+Acc.
+Models of cases 11-$11$
+in Table 3
+9 rewrite_input_dtype
+Output name
+used multiple times.
+Crash GPT-2 (op. 10)
+10
+eliminate_consecutive_idempotent_ops
+Incompatible shape
+error.
+Crash
+Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12),
+R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
+11
+eliminate_duplicate_initializer
+Erase output
+assertion failed.
+Crash
+Faster R-CNN: R-50-FPN-qdq (12)
+12 fuse_bn_into_conv
+Duplicate initializer
+introduced.
+Crash
+RetinaNet (op. 9)
+13
+eliminate_shape_gather
+eliminate_nop_reshape
+Added initializers
+not used by the model.
+Warn.
+Faster R-CNN: R-50-FPN-int8 (op. 12)
+14 All Passes Invalid input scale. Crash
+VGG-16-BN (op. 7), VGG-19-BN (op. 7),
+ResNet101_DUC_HDC (op. 7)
+15 fuse_bn_into_conv Bad graph structure. Acc.
+Efficient-Lite4 (op. 11),
+SSD (op. 12)
+6.2. Output Label Differences in Optimized Classification & Object Detection Models
 Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models.
@@
 Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected.
-Among the $93$ classification models tested, $28$ ($30.1$%) exhibited discrepancies in output labels.
-Similarly, $5$ out of $30$ ($16.6$%) object detection models presented output differences.
-The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($<0.001\%$) while all others showing identical results across the original and the optimized model versions.
-The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction.
-Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues.
-Our methodology identified these inconsistencies, which have been reported to the ONNX Optimizer developers.
-Figure 2: Illustrative example of problematic label classification and object detection between original unoptimized (blue) and optimized (orange) models.
-The bounding box location is affected for both bird instances, while the object label prediction is affected by model optimization on just the right bird instance, changing the label from bird to jet.
-TABLE II: Output label accuracy differences across classification (CL) and object detection (OD) models, showing the type, opset versions, and accuracy differences for top-1, top-5, and top-10 output predictions, between the original and optimized (using the default ONNX Optimizer setting) versions of each model, along with the issue related to the observed differences.
-# Model Type Opset(s) Top 1 Top 5 Top 10 Issue
-1 AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
-2 DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
-3 GoogleNet [62] CL 7-9 0% 0% 0.04% Format Version Change
-4 Inception-1 [61] CL 7-9 0% 0% 0.25% Format Version Change
-5 Inception-2 [61] CL 7-9 0% 0.02% 0.05% Format Version Change
-6 MobileNet v2 [24, 56] CL 7-9 0% 0% 0.02% Format Version Change
-7 ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
-8 ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
-9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
-10 ShuffleNet-v1 [74] CL 7-9 0% 0.7% 1% Format Version Change
-11 SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
-12 EfficientNet-Lite4 [63] CL 11 0% 0.04% 1% Graph Malformation
-13 SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
+Among the $93$ classification models tested, $28$ (30.130.1%) exhibited discrepancies in output labels.
+Similarly, $5$ out of $30$ (16.616.6%) object detection models presented output differences.
+The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($0.001\%$) while all others showing identical results across the original and the optimized model versions.
+The differences impacted the top-5 and top-10 predictions, while top-1 predictions remained unchanged.
+Although top-1 accuracy is often emphasized, variations in lower-ranked predictions can also reveal important issues.
+Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10).
+Using OODTE, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
+Table 3.
+Output label accuracy differences across classification (CL) and object detection (OD) models, showing type, opset version (Ver.), and accuracy gaps for top-1 (T-1), top-5 (T-5), and top-10 (T-10) predictions between original and ONNX-optimized models, along with the related issue (Is.): Format Version (FV) or Malformed Graph (MG).
+# Model Type Ver. T-1 T-5 T-10 Is.
+1 AlexNet (alexnet, ) CL 7-9 0% 0% 0.02% FV
+2 DenseNet121 (densenet, ) CL 7-9 0% 0% 0.02% FV
+3 GoogleNet (googlenet, ) CL 7-9 0% 0% 0.04% FV
+4 Inception-1 (inception, ) CL 7-9 0% 0% 0.25% FV
+5 Inception-2 (inception, ) CL 7-9 0% 0.02% 0.05% FV
+6 MobileNet v2 (mobilenet, ; mobilenetv2, ) CL 7-9 0% 0% 0.02% FV
+7 ResNet18 (resnet, ) CL 7 0% 0.02% 0.02% FV
+8 ResNet34-v2 (resnetv2, ) CL 7 0% 0.02% 0.02% FV
+9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
+10 ShuffleNet-v1 (shufflenet, ) CL 7-9 0% 0.7% 1% FV
+11 SqueezeNet 1.0 (squeezenet, ) CL 7-9 0% 0% 0.02% FV
+12 EfficientNet-Lite4 (efficientnet, ) CL 11 0% 0.04% 1% MG
+13 SSD (ssd, ) OD 12 0% 0% 0.04% MG
 Overall, we detected two primary issues related to output label differences between the original and optimized models:
-Implicit ONNX format version upgrade:
+Implicit ONNX format version upgrade::
 The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication.
 This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1).
-Despite being outdated, these opset versions are still prevalent, representing approximately 41% of the models in the ONNX Model Hub.
+Despite being outdated, these opset versions are still prevalent, representing approximately 4141% of the models in the ONNX Model Hub (presented in Table 3, with the issue labeled as “FV”, indicating a problem with the format version.).
 While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions.
 We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
-Problematic fusion pass:
-The pass fuse_bn_to_conv fuses Batch Normalization and Convolution operators within the model.
-In particular, the pass attempts to fuse BatchNormalization nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node.
+Problematic fusion pass::
+The pass fuse_bn_into_conv
+fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNorma-
+lization
+nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node.
 We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model.
@@
 This behavior is illustrated on the right side of Figure 2.
-This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions.
+This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions (presented in Table 3, with the issue labeled as “MG”, indicating a malformed graph).
 Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided.
 We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
-Table II shows label differences in top-K outputs for classification and object detection models, highlighting the causes.
-As an example, Figure 2 illustrates how an optimized model (orange) can misclassify and alter bounding boxes compared to the original model (blue).
-Figure 3: Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
+Table 3 shows label differences in top-K outputs for classification and object detection models, highlighting the causes.
+Table 4.
+Object detection metric differences between the original and optimized versions, using the default ONNX Optimizer setting.
+For each model, we present the average F1 score (F1), precision (Prec.), and recall (Rec.), as well as mAP and IoU.
+Model Op. F1 Prec. Rec. mAP IoU
+SSD (ssd, ) 12 99.99% 99.99% 99.99% 99.99% 100%
+Tiny
+YOLOv3 (Redmon2018YOLOv3AI, )
+11 99.18% 99.18% 99.18% 99.99% 89.35%
+YOLOv2 (yolov2, ) 9 99.95% 99.95% 99.95% 99.94% 88.49%
+YOLOv3 (Redmon2018YOLOv3AI, ) 10 99.45% 99.45% 99.45% 99.99% 84.81%
+YOLOv3 (Redmon2018YOLOv3AI, ) 12 99.45% 99.45% 99.45% 99.99% 84.81%
+Figure 2.
+Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
 The nodes affected are highlighted.
-The result should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
-TABLE III: Object detection metric differences across the original and optimized versions of object detection models, using the default ONNX Optimizer setting.
-The table presets the name, opset, average F1 score, precision, recall, mean Average Precision (MAP) and average Intersection over Union (IoU scores), for different IoU thresholds ($50$%, $75$%, and $90$%).
-Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
-SSD [39] 12 95.28% 95.28% 95.28% 86.62% 97.37%
-Tiny YOLOv3 [53] 11 99.56% 99.56% 99.56% 90.57% 99.61%
-YOLOv2 [52] 9 99.95% 99.95% 99.95% 90.85% 88.50%
-YOLOv3 [53] 10 98.83% 98.83% 98.83% 86.66% 98.78%
-YOLOv3 [53] 12 98.83% 98.83% 98.83% 86.65% 98.78%
-(a) IoU threshold of 50%.
-Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
-SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
-Tiny YOLOv3 [53] 11 99.36% 99.36% 99.36% 89.97% 99.61%
-YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
-YOLOv3 [53] 10 97.87% 97.87% 97.87% 85.71% 98.78%
-YOLOv3 [53] 12 97.87% 97.87% 97.87% 85.71% 98.78%
-(b) IoU threshold of 75%.
-Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
-SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
-Tiny YOLOv3 [53] 11 99.34% 99.34% 99.34% 89.97% 99.61%
-YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
-YOLOv3 [53] 10 97.8% 97.8% 97.8% 85.69% 98.78%
-YOLOv3 [53] 12 97.8% 97.8% 97.8% 85.69% 98.78%
-(c) IoU threshold of 90%.
-VI-C Bounding Box Differences in Object Detection Models
-We observed notable differences in bounding box outputs between the original and optimized versions of object detection models.
-The results are presented in detail in Table III with the following main findings:
-YOLOv3 Models:
-Both YOLOv3 models (op.set $10$ and $12$) achieve high average IoU, F1 score, precision, and recall—all above $98$%.
-Although the optimized model occasionally misplaces bounding boxes, these errors do not cause misclassifications, as none were found in the top-10 label predictions.
-YOLOv2 Model:
-YOLOv2 shows lower IoU values, likely because it uses multiple bounding boxes per detected object.
-Small differences between the original and optimized models accumulate, lowering IoU.
-However, high F1, precision, and recall indicate that detection performance remains largely unaffected, as demonstrated by the bounding boxes corresponding to the bird on the left in Figure 2.
-SSD Model:
-In contrast, SSD exhibits a $4.8$% drop in F1, precision, and recall between original and optimized models, resulting in 2.92.92.92.96666% misclassification rates despite maintaining a high IoU ($>97$%).
-This behavior is illustrated in Figure 2 by the bounding boxes associated with the bird on the right.
-All models show a substantial mAP difference of up to $14.31$% between original and optimized model versions.
-This suggests the optimizer may cause certain classes to be missed, affecting both object detection and semantic segmentation.
-Consequently, even if detected objects are classified correctly, some objects in the scene might be overlooked, which could significantly impact real-world applications.
-VI-D Optimizer Warnings
-We observed two issues related to the addition of redundant elements, specifically unused initializer entries introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes.
+It should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
+Bounding Box Differences in Object Detection Models We observed some differences in the bounding box outputs between the original and optimized versions of object detection models.
+The results are presented in detail in Table 4 with the following main findings: 1.
+Both YOLOv3 models (opset $10$ and $12$) achieve high average IoU, F1 score, precision, and recall, all above 99.4599.45%.
+However, the optimized model occasionally misplaces bounding boxes—resulting in IoU drops of up to 15.1915.19%, these differences do not lead to misclassifications, as none were observed in the top ranked label predictions.
+A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall).
+6.3. Optimizer Warnings
+We observed two issues related to the addition of redundant elements (unused initializer entries), introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes.
 When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’.
@@
 Upon manual inspection, we found no significant differences in model size.
-However, this issue warrants further investigation for potential side effects.
+However, this issue warrants further investigation for potential side effects towards model performance.
 We have reported it to the optimizer developers for additional review.
-VII Tool Usability
+7. Tool Usability
 OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness.
@@
 This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added.
-By default, OODTE loads models from the official ONNX Model Hub [7], which was used in our experiments.
-However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [6], it can fetch and use models without further modification.
+By default, OODTE loads models from the official ONNX Model Hub (onnxmodelhub, ), which was used in our experiments.
+However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories (onnxhubapi, ), it can fetch and use models without further modification.
 The same flexibility applies to datasets.
-OODTE can fetch datasets using packages like the datasets [3] package or from local storage.
+OODTE can fetch datasets using packages like the datasets (datasets, ) package or from local storage.
 Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
-VIII Threats To Validity
-We list the external and internal threats to validity in our evaluation below.
-VIII-A External Threats
-There are three primary external threats to the validity of our findings.
-(1) Model Selection: Our evaluation included 130 models sourced from the official ONNX Model Hub, covering widely used categories such as classification, object detection, and text comprehension and generation.
-These models were selected for their maturity, accessibility, and broad adoption across academic and industrial domains.
-While we aimed to create a representative and diverse experimental set, the selection is not exhaustive.
-Models with opset versions below 7 were excluded due to incompatibility with the current ONNX and ONNX Runtime versions, which may limit our findings to more recent models.
-Additionally, to keep the evaluation tractable, we did not include certain model types—such as diffusion models and reinforcement learning—which may behave differently under optimization.
-Although our tool supports any ONNX-compliant model hub, the exclusion of these model types presents a potential threat to the generalizability of our results.
-We plan to address this limitation in future work.
-To keep the experiments tractable, we did not evaluate model types such as diffusion models and reinforcement learning in this work; however, we plan to include them in future studies.
-(2) Dataset Selection: Our experimental evaluation relies on widely used, real-world benchmark datasets tailored to the specific tasks of the models under test.
-While this approach ensures relevance and practical applicability—e.g., using ILSVRC [55] for image classification and object detection models like ResNet [22] and GoogleNet [62], and leveraging established HuggingFace datasets [3] for NLP tasks—it may limit generalizability.
-The use of popular datasets could introduce bias toward well-studied scenarios and may not reflect performance on less common or emerging tasks.
-(3) Devices Used: Our experiments were conducted on two hardware platforms: an Intel i5 and an Intel i7 processor.
-While we verified consistency by (a) randomly sampling models of each type and running them on both devices with consistent results, and (b) referencing prior work by Louloudakis et al.
-[41], which found that optimization does not affect accuracy across processing devices, this still presents a potential threat to validity.
-The limited range of hardware may not capture performance or accuracy variations that could arise on other architectures, such as GPUs or edge devices, and may limit the generalizability of our findings.
-VIII-B Internal Threats
-We identify the following internal threat to validity in our findings.
-Optimization Pass Ordering: The order in which optimization passes are applied can significantly affect model performance and behavior, as highlighted in prior work [31].
-This is a well-known challenge in compiler research.
-While this work does not explore the impact of pass ordering in depth, we adopted two evaluation modes in OODTE: (1) the default ONNX Optimizer configuration, which includes commonly used passes such as fuse and eliminate, and (2) a configuration where all 47 available optimization passes were applied atomically.
-Although this strategy does not account for potential issues arising from specific pass combinations or orderings, it has proven effective in uncovering correctness issues in practice.
+8. Threats To Validity
+We list the threats to validity in our evaluation below:
+Model Selection: Our evaluation covered 130 models from the official ONNX Model Hub, spanning classification, object detection, as well as text comprehension and generation.
+Models were chosen for their maturity, accessibility, and broad adoption.
+aimed to create a representative and diverse experimental set, although the selection is not exhaustive.
+We also excluded models with opset versions below 7 due to incompatibility with current ONNX and ONNX Runtime versions, limiting our results to newer models.
+To keep the study tractable, diffusion and reinforcement learning models were omitted; we plan to include these in future work.
+Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (onnx, ) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (onnxsimplifier, )).
+Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies.
+While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing.
+For example, OODTE could support OpenVINO (openvino, ) and Apache TVM (tvm, ) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
+Optimization Pass Ordering:
+The order of optimization passes can significantly affect model performance and behavior (OptimizationOrder, ), a well-known challenge in compiler research.
+While we do not study pass ordering in depth, OODTE adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically.
+Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice.
 Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
-IX Discussion
+9. Discussion
 We summarize below our findings for the different model types in our experiments.
+(1) Optimizations can affect label accuracy in classification models.
+We identified two primary causes for label accuracy differences after optimization: (a) Legacy opset upgrade: Some models used older ONNX opsets, and the optimizer implicitly and improperly upgraded their format versions.
+Classification models built on these legacy opsets were particularly affected.
+(b) Problematic graph node generation: The fuse_bn_into_conv pass, which fuses batch normalization into convolutional layers, can unintentionally alter model behavior and particularly affects classification models that rely on convolutions.
+Although text generation and comprehension models are also large, manual inspection showed that the optimizer applies only a few passes to them.
+Most optimizations are skipped because the targeted layer patterns (e.g., convolutions) are uncommon in transformer-based architectures used for language tasks.
 (2) Object detection models crash more frequently when optimized.
@@
 This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
-While text generation and comprehension models are also typically large, manual inspection revealed that the ONNX Optimizer applies only a limited set of optimizations to them.
-Most passes are skipped, as many of the layer patterns targeted by the optimizer (e.g., convolutions)—common in classification and object detection models—are not typically present in transformer-based architectures used for language tasks.
-(2) Optimization can reduce the accuracy of object detection models.
-While some object detection models remained relatively stable after optimization, others showed notable performance degradation.
-For example, the YOLOv3 model maintained high detection accuracy with no misclassifications, although it exhibited minor shifts in bounding box positions.
-YOLOv2 also showed stable performance, with slightly lower Intersection over Union (IoU) due to its use of multiple bounding boxes per object, but still achieved strong detection metrics overall.
-In contrast, the SSD model experienced a significant decline in F1 score, precision, and recall, leading to higher misclassification rates despite retaining a high IoU.
-Across all models, we observed substantial drops in mean Average Precision (mAP), with reductions of up to 14.31%, suggesting that optimization may cause certain object classes to be missed.
-This raises concerns about the reliability of optimized models in object detection and segmentation tasks.
-(3) Optimizations can affect label accuracy in classification models.
-We identified two primary causes for differences in label accuracy following optimization.
-(a) The first relates to the implicit and improper upgrade of the ONNX format version in models using older opset versions.
-A significant number of classification models were built on these legacy opsets, making them particularly susceptible to this issue.
-(b) The second involves the generation of problematic graph nodes, particularly in relation to the fuse_bn_into_conv pass.
-This pass, which fuses batch normalization into convolutional layers, has a strong impact on classification models and can lead to unintended changes in model behavior after optimization.
+(3) Optimization can affect bounding boxes of object detection models.
+Regarding accuracy, object detection models remained largely stable.
+We observed only a negligible decrease in the accuracy of SSD on top-10 accuracy predictions (0.040.04%), while the other models showed no difference between the original and optimized versions.
+However, the optimized models occasionally misaligned bounding boxes, leading to drops in IoU value of up to 15.1915.19%.
 (4) Outputs from text comprehension and generation models were not affected by optimization.
-From the $8$ model instances under test related to text comprehension and generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model.
+Among the 8 text comprehension and generation models tested, we observed only one crash (in GPT-2) and a small number of samples (8) with BLEU-score differences between the original and optimized models for the top-3 token selections.
 Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others.
 OODTE was able to successfully isolate the optimization passes responsible for the output differences for different model types.
-X Conclusion
+10. Conclusion
 We propose OODTE, a tool for the differential testing of the ONNX Optimizer.
 Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks.
-We detected $9.2$% of model instances that crashed when using the primary optimizer strategies, while $30$% of classification models and $16.6$% of object detection models showed differences between the source and optimized versions, while text generation models remained robust.
-Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
-XI Availability
-For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.
+We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust.
+Additionally, OODTE identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.

### v5

@@
-OODTE: A Differential Testing Engine for the ONNX Optimizer
+DiTOX: Fault Detection and Localization in the ONNX Optimizer
 Abstract.
-With over 760 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models.
-Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated.
-In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer.
-OODTE adopts a straightforward yet powerful differential testing, fault localization, and evaluation methodology, which can be readily adapted for use with other compiler optimizers.
-Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization.
-When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity.
-We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis.
-Our evaluation revealed that 9.29.2% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies.
-Additionally, 3030% of classification models and 16.616.6% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization.
-In total, OODTE uncovered 15 issues, 14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly.
-All findings were reported to the ONNX Optimizer development team.
-OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
+The ONNX Optimizer, part of the official ONNX repository and widely adopted for graph-level model optimizations, is used by default to optimize ONNX models.
+Despite its popularity, its ability to preserve model correctness has not been systematically evaluated.
+We present DiTOX, an automated framework for comprehensively assessing the correctness of the ONNX Optimizer using differential testing, fault localization, and evaluation techniques that generalize to other compiler optimizers.
+DiTOX applies optimization passes to a corpus of ONNX models, executes both original and optimized versions on user-defined inputs, and detects discrepancies in behavior or optimizer failures.
+When divergences are observed, DiTOX isolates the responsible optimization pass through iterative, fine-grained analysis.
+We evaluated DiTOX on 130 models from the ONNX Model Hub spanning vision and language tasks.
+We found that 9.2% of model instances crashed the optimizer or produced invalid models under default settings.
+Moreover, output discrepancies occurred in 30% of classification models and 16.6% of object detection and segmentation models, while text-based models were largely robust.
+Overall, DiTOX uncovered 15 issues—14 previously unknown—affecting 9 of the 47 optimization passes as well as the optimizer infrastructure.
+All issues were reported to the ONNX Optimizer developers.
+Our results demonstrate that DiTOX provides a simple and effective approach for validating AI model optimizers and is readily extensible beyond ONNX.
+Software Engineering, Software Testing, Fault Localization, Differential Testing, ONNX Optimizer
+conference: Proceedings of the 35th ACM SIGPLAN International Conference on Compiler Construction; January 31 – February 1, 2026; Sydney, NSW, Australiabooktitle: Proceedings of the 35th ACM SIGPLAN International Conference on Compiler Construction ($C$ ’26), January 31 – February 1, 2026, Sydney, NSW, Australiaccs: Software and its engineering Empirical software validationccs: Software and its engineering Software defect analysisccs: Software and its engineering Operational analysis
 1. Introduction
@@
 As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies.
-To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (speedisallyouneed, ; CompilerGym, ).
-Widely-used AI compilers and frameworks, such as Apache TVM (tvm, ) and MLIR (lattner2020mlircompilerinfrastructureend, ), incorporate their own native sets of optimization passes for efficient use of hardware.
+To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (Chen et al., 2023; Cummins et al., 2022).
+Widely-used AI compilers and frameworks, such as Apache TVM (Chen et al., 2018) and MLIR (Lattner et al., 2020a), incorporate their own native sets of optimization passes for efficient use of hardware.
 These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
-Models can also be represented using industry-standard formats, such as ONNX (onnxsite, ).
+Models can also be represented using industry-standard formats, such as ONNX (43).
 By applying transformations to models in these formats, it is possible to reduce model size and improve performance.
-For ONNX in particular, the ONNX Optimizer (onnxoptimizer, ) serves as the default tool for performing graph-level optimizations.
+For ONNX in particular, the ONNX Optimizer (40) serves as the default tool for performing graph-level optimizations.
 As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy.
@@
 We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models.
-Ideally, such transformations should not introduce any deviations in model behavior.
-To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a suite that detects and localizes faults in the ONNX Optimizer.
+To this end, we propose the Differential Testing of the ONNX Optimizer (DiTOX) [^1] The source code of DiTOX, along with usage instructions, is available at https://github.com/luludak/DiTOX.
+(Louloudakis, 2025), a suite that detects and localizes faults in the ONNX Optimizer.
 Problem Being Solved
-OODTE solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations.
-In particular, OODTE employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
+DiTOX solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations.
+In particular, DiTOX employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
 Case Studies
-To evaluate the effectiveness of OODTE, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation.
-All models were obtained from the official ONNX Model Hub (onnxmodelhub, ).
+To evaluate the effectiveness of DiTOX, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation.
+All models were obtained from the official ONNX Model Hub (39).
 Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization.
 When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
-We focus on the ONNX Optimizer due to its impact and popularity (over 760 GitHub stars) as a model optimization tool.
-In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (onnxsimplifier, ), which has received over 4,000 GitHub stars.
+We focus on the ONNX Optimizer due to its impact and popularity (over 770 GitHub stars) as a model optimization tool.
+In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (42), which has received over 4,000 GitHub stars.
 However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes.
-With OODTE, we aim to enhance the effectiveness of fault localization and detection in the optimizer.
+With DiTOX, we aim to enhance the effectiveness of fault localization and detection in the optimizer.
 However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
@@
 All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution.
-OODTE effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer.
+DiTOX effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer.
 These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
@@
 Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters.
-These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (softmax, ).
+These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (Pearce et al., 2021).
 2.1.2. Object Detection and Semantic Segmentation
@@
 These models handle tasks such as summarization, recognition, and sentiment analysis.
-Text generation models, especially transformers (attention, ), use contextual representations to iteratively predict the next token using a sliding window approach.
+Text generation models, especially transformers (Vaswani et al., 2017), use contextual representations to iteratively predict the next token using a sliding window approach.
 This generative process continues until a stopping condition is met.
@@
 This may involve techniques such as node pruning, node fusion, and constant folding, among others.
-Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (tvm, ; lattner2020mlircompilerinfrastructureend, ) incorporating such techniques.
+Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (Chen et al., 2018; Lattner et al., 2020a) incorporating such techniques.
 Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
@@
 The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes.
-It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency.
+It consists of a set of $47$ passes [^2] The complete list of passes, each accompanied by a brief description, can be found at https://github.com/luludak/DiTOX/blob/master/passes.txt.
+grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency.
 Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead.
@@
 Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types.
-In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_pre-
-dict
-, which isolate the initializer and inference subgraphs, respectively.
+In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_predict, which isolate the initializer and inference subgraphs, respectively.
 By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
 2.4. Model Output Comparison Methods
-Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts.
+Based on the model and output type, DiTOX uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts.
 In particular:
 Classification models::
-We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (kendall, ), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
-Object Detection & Semantic Segmentation models::
-Ea-
-ch prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score.
+We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (Kendall, 1938), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
+Object Detection and Semantic Segmentation models::
+Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score.
 For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity.
@@
 Text generation models::
-We convert token outputs to plain text and use the BLEU Score (BLEU, ) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
+We convert token outputs to plain text and use the BLEU Score (Papineni et al., 2002) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
 Binary sentiment analysis models::
@@
 3. System Architecture & Methodology
-Figure 1. Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
-The architecture of OODTE and its modules is shown in Figure 1.
+Figure 1. Architecture of DiTOX with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
+The architecture of DiTOX and its modules is shown in Figure 1.
 Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results.
-If output discrepancies are detected, OODTE iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference.
+If output discrepancies are detected, DiTOX iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference.
 All results are stored in report files for analysis and fault localization.
-OODTE consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
+DiTOX consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
 It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module.
 The modules in Figure 1 are detailed in the following sections.
+DiTOX currently isolates bugs at the level of individual passes, relying on compiler error messages to identify the cause.
+For more effective fault localization, future enhancements can explore data flow analysis and delta debugging within the identified pass to identify the root cause.
 3.1. Model Orchestrator Module
@@
 The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository.
-At the moment, OODTE supports the official ONNX Model Hub (onnxmodelhub, ), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories.
+At the moment, DiTOX supports the official ONNX Model Hub (39), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories.
 In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification.
@@
 It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘).
-OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations.
-This feature is essential throughout the OODTE pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth.
-To identify the pass responsible for problematic behavior, OODTE first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified.
+DiTOX supports the utilization of these passes for testing, in order to explore all optimizer operations.
+This feature is essential throughout the DiTOX pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth.
+To identify the pass responsible for problematic behavior, DiTOX first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified.
 Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
@@
 In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions.
-It then performs model inference, using ONNX Runtime (onnxruntime, ), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component.
-OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews.
-It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (datasets, ) package.
+It then performs model inference, using ONNX Runtime (41), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component.
+DiTOX supports different types of datasets, based on the type of models being examined - from images to IMDB reviews.
+It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (8) package.
 To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts.
@@
 This module performs comparison against the results, by selecting the appropriate comparator based on the model type.
-For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (kendall, ).
-For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (IoUBoundingBox, )), Precision, Recall, and F1 Score (F1Score, ) given an IoU threshold.
-For text generation models (e.g., GPT-2), it utilizes BLEU (BLEU, ) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A).
-In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
+For classification types, DiTOX supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (Kendall, 1938).
+For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (Števuliáková and Hurtik, 2023)), Precision, Recall, and F1 Score (Christen et al., 2023) given an IoU threshold.
+For text generation models (e.g., GPT-2), it utilizes BLEU (Papineni et al., 2002) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A).
+In addition DiTOX is build in a manner that new comparators can be added, in cases that a new category of models is tested.
 3.5. Fault Localization
 Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format.
-If a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module.
+If a difference is detected across the original and the optimized model, then DiTOX retrofits the reported data to the Optimizer Module.
 Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.
-OODTE applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model.
-In addition, OODTE registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model).
-Upon detecting an issue or a problematic behavior, OODTE attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model.
+DiTOX applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model.
+In addition, DiTOX registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model).
+Upon detecting an issue or a problematic behavior, DiTOX attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model.
 This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
@@
 Table 1.
-Comparison of features between OODTE and related work.
+Comparison of features between DiTOX and related work.
 Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt.
@@
 Tool Technique ON NX? Flt. Loc. Bug Types Gr. Opt. Passes Acc. Cr. Malf. T. Compilers
-OODTE
-Per-pass fault
-localization,
-differential
-testing
-Logic, API versioning,
-graph input/parameters/
-transformations
-Fuse/
-eliminate/
-rewrite
-ONNX
-Optimizer (onnxoptimizer, )
-NNSmith (NNSmith, )
-Grammar-
-based fuzzing
-†,‡ $\times$
-Conversions, semantics,
-gen.transformations,
-graph rewrites
-Conv2D
-inputs
-rewrite
-$\times$ $\times$
-TVM (tvm, ),
-PyTorch JIT (pytorch, ),
-TensorRT (tensorrt, ),
-PT2 (pytorch2, ), ONNX
-Runtime (onnxruntime, ),
-TF XLA (tfxla, ),
-TFLite (tensorflow, )
-MLIRSmith (MLIRSmith, )
-Grammar-
-based fuzzing
-$\times$ $\times$
-Conversions (dialects),
-generic transformations,
-logic (dialects)
-$\times$ $\times$ $\times$ $\times$
-MLIR (MLIR, ),
-LLVM (LLVM, )
-MLIRod (SuoMLIR, )
-Operator
-dependency
-analysis/
-fuzzing
-$\times$ $\times$
-Conversions (dialects),
-gen. transformations
-$\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
-SYNTH-
-FUZZ (synthfuzz, )
-Grammar-
-based fuzzing
-$\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
-HirGen (HirGen, )
-Graph IR
-fuzzing
-$\times$ $\times$
-Numeric computations,
-exceptions, types, memory
-$\times$ $\times$ $\times$ $\times$ TVM (tvm, )
-TVMFuzz (TVMFuzz, )
-Low-level
-IR fuzzing
-$\times$ $\times$
-Numeric computations,
-graph rewrites
-Constant
-folding
-$\times$ $\times$ TVM (tvm, )
-Tzer (TZER, )
-Coverage-
-guided fuzzing
-(Low-level IR)
-$\times$ $\times$
-API inconsistency/misuse,
-memory, types
-$\times$ $\times$ TVM (tvm, )
-GenCoG (GenCoG, )
-Constraint
-solving,
-High-level
-IR Fuzzing
-$\times$ $\times$
-Memory, tensor shape
-/data, operator attributes,
-logic
-Operator
-fusion
-$\times$ $\times$ TVM (tvm, )
-DeepDiffer (DeepDiffer, )
-Low-level IR/
-coverage-
-guided fuzzing,
-differential
-testing
-$\times$ $\times$
-Type, memory, API misuse,
-numeric computations,
-exception handling
-$\times$ $\times$ $\times$ TVM (tvm, )
-WhiteFox (WhiteFox, )
-LM analysis,
-fuzzing
-$\times$ $\times$
-Memory, API misuse,
-concurrency, logic,
-security vulnerabilities
-$\times$ $\times$
-TF XLA (tfxla, ),
-PyTorch
-Inductor (pytorch, ),
-LLVM (LLVM, )
-TorchProbe (Su_2023, )
-Feature-
-based fuzzing
-$\times$ $\times$
-Memory, API misuse,
-logic
-$\times$ $\times$ $\times$
-PyTorch (pytorch, ),
-Triton (TilletTriton, )
-MT-
-DLComp (Xiao_2022, )
-Metamorphic
-testing
-$\times$ $\times$
-Memory, dead code,
-layout optimization,
-numeric accuracy
-Conv2D
-layout
-change
-$\times$ $\times$
-TVM (tvm, ),
-Glow (Glow, ),
-NNFusion (NNFusion, ),
-TF XLA (tfxla, )
-FetaFix (FetaFix, )
-Fault localiza-
-tion,
-differential
-testing
-$\times$
-Preprocessing, input
-dimensions, parameters,
-hyperparameters, tensor
-shape, graph structure
-$\times$ $\times$ $\times$ $\times$
-Converters across
-PyTorch (pytorch, ),
-Keras (chollet2015keras, ),
-TF (tensorflow, ), TFLite (tensorflow, )
-In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (DeepDiffer, )) or LLM analysis (e.g., WhiteFox (WhiteFox, )).
+DiTOX Per-pass fault localization, differential testing Logic, API versioning, graph input/parameters/ transformations Fuse/ eliminate/ rewrite ONNX Optimizer (40)
+NNSmith (Liu et al., 2023) Grammar- based fuzzing †,‡ $\times$ Conversions, semantics, transformations, graph rewrites Conv2D inputs rewrite $\times$ $\times$ TVM (Chen et al., 2018), PyTorch JIT (Paszke et al., 2019), TensorRT (NVIDIA, 2019), PT2 (Ansel et al., 2024), ONNX Runtime (41), TF XLA (Sabne, 2020), TFLite (Abadi et al., 2016)
+MLIRSmith (Wang et al., 2023a) Grammar- based fuzzing $\times$ $\times$ Conversions (dialects), generic transformations, logic (dialects) $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b), LLVM (Lattner and Adve, 2004)
+MLIRod (Suo et al., 2024) Operator dependency/ fuzzing $\times$ $\times$ Conversions (dialects), transformations $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b)
+SYNTH- FUZZ (Limpanukorn et al., 2025) Grammar- based fuzzing $\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (Lattner et al., 2020b)
+HirGen (Ma et al., 2023) Graph IR fuzzing $\times$ $\times$ Numeric computations, exceptions, types, memory $\times$ $\times$ $\times$ $\times$ TVM (Chen et al., 2018)
+TVMFuzz (64) Low-level IR fuzzing $\times$ $\times$ Numeric computations, graph rewrites Constant folding $\times$ $\times$ TVM (Chen et al., 2018)
+Tzer (Liu et al., 2022) Coverage- guided fuzzing (Low-level IR) $\times$ $\times$ API inconsistency/misuse, memory, types $\times$ $\times$ TVM (Chen et al., 2018)
+GenCoG (Wang et al., 2023b) Constraint solving, High-level IR Fuzzing $\times$ $\times$ Memory, tensor shape /data, operator attributes, logic Operator fusion $\times$ $\times$ TVM (Chen et al., 2018)
+DeepDiffer (Lin et al., 2023) Low-level IR/ coverage- guided fuzzing, differential testing $\times$ $\times$ Type, memory, API misuse, numeric computations, exception handling $\times$ $\times$ $\times$ TVM (Chen et al., 2018)
+WhiteFox (Yang et al., 2024) LM analysis, fuzzing $\times$ $\times$ Memory, API misuse, concurrency, logic, security vulnerabilities $\times$ $\times$ TF XLA (Sabne, 2020), PyTorch Inductor (Paszke et al., 2019), LLVM (Lattner and Adve, 2004)
+TorchProbe (Su et al., 2023) Feature- based fuzzing $\times$ $\times$ Memory, API misuse, logic $\times$ $\times$ $\times$ PyTorch (Paszke et al., 2019), Triton (Tillet et al., 2019)
+MT- DLComp (Xiao et al., 2022) Metamorphic testing $\times$ $\times$ Memory, dead code, layout optimization, numeric accuracy Conv2D layout change $\times$ $\times$ TVM (Chen et al., 2018), Glow (Rotem et al., 2019), NNFusion (Ma et al., 2020), TF XLA (Sabne, 2020)
+FetaFix (Louloudakis et al., 2025) Fault localiza- tion, differential testing $\times$ Preprocessing, input dimensions, parameters, hyperparameters, tensor shape, graph structure $\times$ $\times$ $\times$ $\times$ Converters across PyTorch (Paszke et al., 2019), Keras (Chollet, 2015), TF (Abadi et al., 2016), TFLite (Abadi et al., 2016)
+In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (Lin et al., 2023)) or LLM analysis (e.g., WhiteFox (Yang et al., 2024)).
 Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization.
-Among these works, only FetaFix (FetaFix, ) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process.
-In contrast, OODTE systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
+Among these works, only FetaFix (Louloudakis et al., 2025) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process.
+In contrast, DiTOX systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
 In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs.
-For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (NNSmith, )) to expose inconsistencies in model inference.
-In contrast, for label inconsistency bugs that OODTE detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (ILSVRC17, ), IMDB (imdbsentiment, )) and model architectures.
-Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (TVMFuzz, ; WhiteFox, ).
-The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (TZER, ; DeepDiffer, ) using empirical measurements to evaluate performance-related faults.
-OODTE currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries), while the integration of precise execution time measurements is left as subject for future work.
-Regarding support for ONNX in Table 1, OODTE is the only tool capable of localizing faults in the ONNX Optimizer.
-Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (WhiteFox, ), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
+For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (Liu et al., 2023)) to expose inconsistencies in model inference.
+In contrast, for label inconsistency bugs that DiTOX detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (Russakovsky et al., 2015), IMDB (Tripathi et al., 2020)) and model architectures.
+Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (64; C.
+Yang, Y.
+Deng, R.
+Lu, J.
+Yao, J.
+Liu, R.
+Jabbarvand, and L.
+Zhang (2024)).
+The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (Liu et al., 2022; Lin et al., 2023) using empirical measurements to evaluate performance-related faults.
+DiTOX currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries).
+Regarding support for ONNX in Table 1, DiTOX is the only tool capable of localizing faults in the ONNX Optimizer.
+Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (Yang et al., 2024), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
 Existing work on bug detection within graph optimization passes (shown as Gr.
@@
 Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations.
-In contrast, OODTE addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
+In contrast, DiTOX addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
 5. Experiment Setup
@@
 When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue.
-OODTE also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
+DiTOX also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
 5.2. DNN Models
-For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (onnxmodelhub, ).
-In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet (mobilenet, ), EfficientNet-Lite4 (efficientnet, ) DenseNet121 (densenet, ), and InceptionV2 (inception, )), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (retinanet, ), Mask/Faster R-CNN R-50-FPN (mask-rcnn, ; faster-rcnn, ), and SSD (ssd, )), and $7$ transformer models (attention, ) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (BERT, )), sentiment analysis (RoBERTa (liu2020roberta, )), text summarization (T5 (T5Paper, )) and generic-purpose text generation (GPT-2 (gpt2, )), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks.
+For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (39).333.
+The complete set of models used in our experiments, along with the full set of our results, is available in the results folder of the DiTOX repository at https://github.com/luludak/DiTOX/blob/master/results/README.md.
+In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNetV2 (Howard et al., 2017), EfficientNet-Lite4 (10) DenseNet121 (Huang et al., 2018), as well as InceptionV2 (Szegedy et al., 2014b)), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (Lin et al., 2020), Mask/Faster R-CNN R-50-FPN (He et al., 2020; Ren et al., 2015), and SSD (Liu et al., 2016)), and $7$ transformer models (Vaswani et al., 2017) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (Devlin et al., 2019)), sentiment analysis (RoBERTa (Liu et al., 2020)), text summarization (T5
+ (Raffel et al., 2020)) and generic-purpose text generation (GPT-2 (21)), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks.
 We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API).
@@
 We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning.
-Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (onnxruntime, ), or an AI compiler such as Apache TVM (tvm, )).
+Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (41), or an AI compiler such as Apache TVM (Chen et al., 2018)).
 5.4. Datasets
 We selected datasets appropriate to each model task.
-For classification and object detection, we used the ILSVRC2017 (ILSVRC17, ) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (resnet, ) and BERT (BERT, )).
-For text comprehension and generation, we used SQuAD-v2 (rajpurkar2018knowdontknowunanswerable, ), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
-Additionally, for summarization and sentiment analysis, we used the IMDB dataset (moviedataset, ; imdbsentiment, ), containing $25000$ entries.
+For classification and object detection, we used the ILSVRC2017 (Russakovsky et al., 2015) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (He et al., 2015) and BERT (Devlin et al., 2019)).
+For text comprehension and generation, we used SQuAD-v2 (Rajpurkar et al., 2018), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior.
+Additionally, for summarization and sentiment analysis, we used the IMDB dataset (Maas et al., 2011; Tripathi et al., 2020), containing $25000$ entries.
 To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
-  1. (1)
-    For image recognition models performing classification tasks, we utilized the Kendall Rank Correlation Coefficient (kendall, ) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$.
-If the KT value was found to be <1<1 for each input, we counted this input as different.
-  2. (2)
+1.
+For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient (Kendall, 1938) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$.
+If the KT value was found to be <1<1, we treated this input as different.
+2.
 For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall.
 We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
-  3. (3)
-    For text generation models performing question answering and summarization, we utilized BLEU score (BLEU, ) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
-  4. (4)
-    For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
+3.
+For text generation models performing question answering and summarization, we utilized BLEU score (Papineni et al., 2002) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
+4. For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
+5.
+To distinguish actual faults from known optimizer limitations, we relied on ONNX Optimizer documentation and accounted for known limitations (e.g., split_init and split_predict are known to not fully comply with the ONNX standard format).
+Aside from issues from these two passes, all other detected behaviors were unexpected and thus classified as faults.
 6. Results
-Following the execution of the full set of experiments.
+Following the execution of the full set of experiments11footnotemark: 1, DiTOX successfully identified $15$ bugs, reported across $11$ separate GitHub issues (11) [^4] An overview of all detected issues using DiTOX can be found at https://github.com/luludak/DiTOX/blob/master/results/detected_issues.txt..
 A complete list of these issues, along with the associated models, is shown in Table 2.
@@
 By default, the optimizer prioritizes passes related to fuse and eliminate operations.
-As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy.
-Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage.
+As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the core (default) strategy of the optimizer.
+Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the overall usage of the optimizer.
 We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
@@
 In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes.
-In particular, R-CNN models, such as Faster R-CNN R-50-FPN (faster-rcnn, ), RetinaNet (retinanet, ), VGG-16, VGG-19 (vgg, ), as well as YOLOv3 (Redmon2018YOLOv3AI, ), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations.
+In particular, R-CNN models, such as Faster R-CNN R-50-FPN (Ren et al., 2015), RetinaNet (Lin et al., 2020), VGG-16, VGG-19 (Simonyan and Zisserman, 2015), as well as YOLOv3 (Redmon and Farhadi, 2018), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations.
 In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash.
@@
 Table 2.
-Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
+Overview of the issues identified using DiTOX, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
+The (Run) label in crashes indicates that the model was successfully optimized but failed during execution.
 # Pass Error Type Models
-1 lift_lexical_references
-Removes used
-value references.
-Crash
-EfficientNet-Lite4 (op. 11), SSD (op. 12)
-2 split_init
-List index
-out of range.
-Crash YOLOv3 (op. 10, 12)
-3
-split_init
-split_predict
-Required Field
-is missing.
-Crash
-DenseNet121 (op. 7-9), YOLOv2 (op. 10),
-SSD (op. 12), Tiny-YOLOv3 (op. 11),
-YOLOv3 (op. 10, 12)
-4 split_predict
-Required inputs
-are missing.
-Crash
-SSD (op. 12), DenseNet121 (op. 7),
-InceptionV2 (op. 6-9)
-5 split_predict Bad spec for node. Crash
-YOLOv3 (op. 10), DenseNet121 (op. 7),
-InceptionV2 (op. 6-9)
-6 split_predict
-Mismatch in number
-of input elements
-Crash
-FCN ResNet-50-int8 (op. 12)
-7 eliminate_shape_gather
-Indices value
-not equal to
-node dims size
-Crash
-Faster R-CNN: R-50-FPN (op. 10)
-8 All Passes
-Implicit ONNX
-format version
-update.
-Acc.
-Models of cases 11-$11$
-in Table 3
-9 rewrite_input_dtype
-Output name
-used multiple times.
-Crash GPT-2 (op. 10)
-10
-eliminate_consecutive_idempotent_ops
-Incompatible shape
-error.
-Crash
-Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12),
-R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
-11
-eliminate_duplicate_initializer
-Erase output
-assertion failed.
-Crash
-Faster R-CNN: R-50-FPN-qdq (12)
-12 fuse_bn_into_conv
-Duplicate initializer
-introduced.
-Crash
-RetinaNet (op. 9)
-13
-eliminate_shape_gather
-eliminate_nop_reshape
-Added initializers
-not used by the model.
-Warn.
-Faster R-CNN: R-50-FPN-int8 (op. 12)
-14 All Passes Invalid input scale. Crash
-VGG-16-BN (op. 7), VGG-19-BN (op. 7),
-ResNet101_DUC_HDC (op. 7)
-15 fuse_bn_into_conv Bad graph structure. Acc.
-Efficient-Lite4 (op. 11),
-SSD (op. 12)
+1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
+2 split_init Out of range List index Crash YOLOv3 (op. 10, 12)
+3 split_init split_predict Missing required field Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11), YOLOv3 (op. 10, 12)
+4 split_predict Missing required inputs Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
+5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
+6 split_predict Mismatch in #input elements Crash FCN ResNet-50-int8 (op. 12)
+7 eliminate_shape_gather Indices value $\neq$ node dims size. Crash Faster R-CNN: R-50-FPN (op. 10)
+8 All Passes Implicit ONNX format version update. Acc. Models of cases 11-$11$ in Table 3
+9 rewrite_input_dtype Multiple usage of output name Crash GPT-2 (op. 10)
+10 eliminate_consecutive _idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
+11 eliminate_duplicate_initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
+12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
+13 eliminate_shape_gather eliminate_nop_reshape Unused initializers Warn. Faster R-CNN: R-50-FPN-int8 (op. 12)
+14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
+15 fuse_bn_into_conv Bad graph structure. Acc. Efficient-Lite4 (op. 11), SSD (op. 12)
 6.2. Output Label Differences in Optimized Classification & Object Detection Models
@@
 Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10).
-Using OODTE, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
+Using DiTOX, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
 Table 3.
@@
 # Model Type Ver. T-1 T-5 T-10 Is.
-1 AlexNet (alexnet, ) CL 7-9 0% 0% 0.02% FV
-2 DenseNet121 (densenet, ) CL 7-9 0% 0% 0.02% FV
-3 GoogleNet (googlenet, ) CL 7-9 0% 0% 0.04% FV
-4 Inception-1 (inception, ) CL 7-9 0% 0% 0.25% FV
-5 Inception-2 (inception, ) CL 7-9 0% 0.02% 0.05% FV
-6 MobileNet v2 (mobilenet, ; mobilenetv2, ) CL 7-9 0% 0% 0.02% FV
-7 ResNet18 (resnet, ) CL 7 0% 0.02% 0.02% FV
-8 ResNet34-v2 (resnetv2, ) CL 7 0% 0.02% 0.02% FV
-9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
-10 ShuffleNet-v1 (shufflenet, ) CL 7-9 0% 0.7% 1% FV
-11 SqueezeNet 1.0 (squeezenet, ) CL 7-9 0% 0% 0.02% FV
-12 EfficientNet-Lite4 (efficientnet, ) CL 11 0% 0.04% 1% MG
-13 SSD (ssd, ) OD 12 0% 0% 0.04% MG
+1 AlexNet (Krizhevsky et al., 2017) CL 7-9 0% 0% 0.02% FV
+2 DenseNet121 (Huang et al., 2018) CL 7-9 0% 0% 0.02% FV
+3 GoogleNet (Szegedy et al., 2014a) CL 7-9 0% 0% 0.04% FV
+4 InceptionV1 (Szegedy et al., 2014b) CL 7-9 0% 0% 0.25% FV
+5 InceptionV2 (Szegedy et al., 2014b) CL 7-9 0% 0.02% 0.05% FV
+6 MobileNetV2 (Howard et al., 2017; Sandler et al., 2018) CL 7-9 0% 0% 0.02% FV
+7 ResNet18 (He et al., 2015) CL 7 0% 0.02% 0.02% FV
+8 ResNet34V2 (He et al., 2016) CL 7 0% 0.02% 0.02% FV
+9 ResNet50-caffe2 (He et al., 2015) CL 7-9 0% 2% 13% FV
+10 ShuffleNetV1 (Zhang et al., 2017) CL 7-9 0% 0.7% 1% FV
+11 SqueezeNet 1.0 (Iandola et al., 2017) CL 7-9 0% 0% 0.02% FV
+12 EfficientNet-Lite4 (10) CL 11 0% 0.04% 1% MG
+13 SSD (Liu et al., 2016) OD 12 0% 0% 0.04% MG
 Overall, we detected two primary issues related to output label differences between the original and optimized models:
@@
 The pass fuse_bn_into_conv
-fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNorma-
-lization
+fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNor-
+malization
 nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node.
@@
 Model Op. F1 Prec. Rec. mAP IoU
-SSD (ssd, ) 12 99.99% 99.99% 99.99% 99.99% 100%
-Tiny
-YOLOv3 (Redmon2018YOLOv3AI, )
-11 99.18% 99.18% 99.18% 99.99% 89.35%
-YOLOv2 (yolov2, ) 9 99.95% 99.95% 99.95% 99.94% 88.49%
-YOLOv3 (Redmon2018YOLOv3AI, ) 10 99.45% 99.45% 99.45% 99.99% 84.81%
-YOLOv3 (Redmon2018YOLOv3AI, ) 12 99.45% 99.45% 99.45% 99.99% 84.81%
+SSD (Liu et al., 2016) 12 99.99% 99.99% 99.99% 99.99% 100%
+Tiny YOLOv3 (Redmon and Farhadi, 2018) 11 99.18% 99.18% 99.18% 99.99% 89.35%
+YOLOv2 (Redmon and Farhadi, 2016) 9 99.95% 99.95% 99.95% 99.94% 88.49%
+YOLOv3 (Redmon and Farhadi, 2018) 10 99.45% 99.45% 99.45% 99.99% 84.81%
+YOLOv3 (Redmon and Farhadi, 2018) 12 99.45% 99.45% 99.45% 99.99% 84.81%
 Figure 2.
@@
 A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall).
+Although this behavior may appear minor, it can lead to undesirable consequences, particularly in safety-critical systems where even small deviations matter.
+While our experiments did not observe bounding box misclassifications, such outcomes cannot be ruled out for other models or different inputs.
 6.3. Optimizer Warnings
@@
 7. Tool Usability
-OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness.
-Currently, OODTE supports differential testing for classification, object detection, and text generation models.
-However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
-In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline.
-This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added.
-By default, OODTE loads models from the official ONNX Model Hub (onnxmodelhub, ), which was used in our experiments.
-However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories (onnxhubapi, ), it can fetch and use models without further modification.
+DiTOX  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness.
+Currently, DiTOX supports differential testing for classification, object detection, and text generation models.
+However, DiTOX can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
+In addition, DiTOX can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline.
+This allows DiTOX to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added.
+By default, DiTOX loads models from the official ONNX Model Hub (39), which was used in our experiments.
+However, by configuring DiTOX to use a custom ONNX Hub that follows the ONNX API specification for model repositories (38), it can fetch and use models without further modification.
 The same flexibility applies to datasets.
-OODTE can fetch datasets using packages like the datasets (datasets, ) package or from local storage.
-Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
+DiTOX can fetch datasets using packages like the datasets (8) package or from local storage.
+8. Tool Generalizability
+Although DiTOX applies differential testing to the ONNX Optimizer, the approach is generic and can be adapted to other AI optimizers or compilers that provide a runtime for model execution and expose optimization passes both collectively and individually through an API.
+Many compiler ecosystems already satisfy these requirements.
+For example, MLIR (Lattner et al., 2020b) offers passes such as graph rewrites, Apache TVM (Chen et al., 2018) provides a Sequential API for composing optimizations like constant folding and subexpression elimination, and OpenVINO (Zunin, 2021) supports a full set of individual passes.
+While we focus on ONNX due to its wide adoption, the methodology of DiTOXis model-format independent and can be applied to other frameworks using their supported model types and runtimes.
+Finally, DiTOX includes comparators for classifiers, object detectors, and transformers; for other model families, users can implement custom comparators by following the already provided comparator API.
 9. Threats To Validity
@@
 To keep the study tractable, diffusion and reinforcement learning models were omitted; we plan to include these in future work.
-Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (onnx, ) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (onnxsimplifier, )).
+Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (43) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (42)).
 Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies.
 While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing.
-For example, OODTE could support OpenVINO (openvino, ) and Apache TVM (tvm, ) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
+For example, DiTOX could support OpenVINO (Zunin, 2021) and Apache TVM (Chen et al., 2018) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
 Optimization Pass Ordering:
-The order of optimization passes can significantly affect model performance and behavior (OptimizationOrder, ), a well-known challenge in compiler research.
-While we do not study pass ordering in depth, OODTE adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically.
+The order of optimization passes can significantly affect model performance and behavior (Kulkarni and Cavazos, 2012), a well-known challenge in compiler research.
+While we do not study pass ordering in depth, DiTOX adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically.
 Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice.
@@
 10. Discussion
-We summarize below our findings for the different model types in our experiments.
+We summarize our findings across different model types below.
 (1) Optimizations can affect label accuracy in classification models.
-We identified two primary causes for label accuracy differences after optimization: (a) Legacy opset upgrade: Some models used older ONNX opsets, and the optimizer implicitly and improperly upgraded their format versions.
-Classification models built on these legacy opsets were particularly affected.
-(b) Problematic graph node generation: The fuse_bn_into_conv pass, which fuses batch normalization into convolutional layers, can unintentionally alter model behavior and particularly affects classification models that rely on convolutions.
-Although text generation and comprehension models are also large, manual inspection showed that the optimizer applies only a few passes to them.
-Most optimizations are skipped because the targeted layer patterns (e.g., convolutions) are uncommon in transformer-based architectures used for language tasks.
-(2) Object detection models crash more frequently when optimized.
-Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models.
-This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
-(3) Optimization can affect bounding boxes of object detection models.
-Regarding accuracy, object detection models remained largely stable.
-We observed only a negligible decrease in the accuracy of SSD on top-10 accuracy predictions (0.040.04%), while the other models showed no difference between the original and optimized versions.
-However, the optimized models occasionally misaligned bounding boxes, leading to drops in IoU value of up to 15.1915.19%.
-(4) Outputs from text comprehension and generation models were not affected by optimization.
-Among the 8 text comprehension and generation models tested, we observed only one crash (in GPT-2) and a small number of samples (8) with BLEU-score differences between the original and optimized models for the top-3 token selections.
-Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others.
-OODTE was able to successfully isolate the optimization passes responsible for the output differences for different model types.
+Label discrepancies in classification models arose primarily from two sources.
+First, models using legacy ONNX opsets were implicitly and incorrectly upgraded by the optimizer, disproportionately affecting classification models.
+Second, the fuse_bn_into_conv pass occasionally generated problematic graph nodes, altering model behavior in convolution-heavy architectures.
+In contrast, text models were largely unaffected, as the optimizer applies few passes to transformer-based architectures where convolutional patterns are rare.
+(2) Object detection models are more prone to optimization crashes.
+Most optimizer crashes occurred in object detection and semantic segmentation models.
+Their larger size and structural complexity trigger more aggressive transformations, increasing exposure to corner cases in optimization passes.
+(3) Optimization can degrade bounding box accuracy.
+While overall detection accuracy remained stable, optimized models occasionally produced misaligned bounding boxes.
+This resulted in IoU drops of up to 15.1915.19%, despite negligible changes in top-10 accuracy (e.g., 0.040.04% for SSD).
+(4) Text models are largely robust to optimization.
+Among the eight text comprehension and generation models, we observed only one optimizer crash (GPT-2) and minor BLEU-score differences affecting eight samples in top-3 token predictions.
+Overall, ONNX optimizations can introduce correctness issues, with susceptibility varying by model type.
+DiTOX effectively isolates the optimization passes responsible for these discrepancies across model classes.
+11. Future Work
+We identify three potential extensions of this work as future research directions:
+(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically on detecting crashes and accuracy faults.
+However, in the broader context of model optimization, it is also important to assess how an optimizer impacts overall model performance.
+In particular, this includes execution time for both CPU and GPU deployments, as well as other factors such as model size.
+DiTOX can be easily extended to support hardware-accelerated model execution, either by leveraging the ONNX Runtime API (41) or by integrating a third-party AI compiler that generates target-specific device code, such as Apache TVM (Chen et al., 2018).
+We leave this direction for future work, as it could provide valuable additional insights.
+(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models.
+However, DiTOX is applicable to a broader range of model types, including speech recognition, super-resolution, and body recognition.
+We leave the exploration of these additional model categories as a direction for future work.
+(3) Examination of the Optimization Pass Order:
+DiTOX supports optimization using a user-defined subset and ordering of passes through configuration.
+Although such experiments could potentially reveal additional faults, we exclude them for tractability, as exhaustively generating, executing, and evaluating all pass permutations is computationally infeasible.
+We leave this direction for future work, potentially in combination with techniques for reducing the permutation search space.
 12. Conclusion
-We propose OODTE, a tool for the differential testing of the ONNX Optimizer.
-Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks.
+We propose DiTOX, a tool for the differential testing of the ONNX Optimizer.
+Using DiTOX, we evaluated $130$ models across classification, object detection, and text generation tasks.
 We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust.
-Additionally, OODTE identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.
+Additionally, DiTOX identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.